### v1

# Why the Agent Made that Decision: Explaining Deep Reinforcement Learning with Vision Masks

Abstract
Due to the inherent lack of transparency in deep neural networks, it is challenging for deep reinforcement learning (DRL) agents to gain trust and acceptance from users, especially in safety-critical applications such as medical diagnosis and military operations. Existing methods for explaining an agent’s decision either require to retrain the agent using models that support explanation generation or rely on perturbation-based techniques to reveal the significance of different input features in the decision making process. However, retraining the agent may compromise its integrity and performance, while perturbation-based methods have limited performance and lack knowledge accumulation or learning capabilities. Moreover, since each perturbation is performed independently, the joint state of the perturbed inputs may not be physically meaningful. To address these challenges, we introduce VisionMask, a standalone explanation model trained end-to-end to identify the most critical regions in the agent’s visual input that can explain its actions. VisionMask is trained in a self-supervised manner without relying on human-generated labels. Importantly, its training does not alter the agent model, hence preserving the agent’s performance and integrity. We evaluate VisionMask on Super Mario Bros (SMB) and three Atari games. Compared to existing methods, VisionMask achieves a 14.9% higher insertion accuracy and a 30.08% higher F1-Score in reproducing original actions from the selected visual explanations. We also present examples illustrating how VisionMask can be used for counterfactual analysis.
1 Introduction
Original State LIME RISE Greydanus et al. SARFA VisionMask
[Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
[Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
Table 1: Visual comparison of saliency maps generated by VisionMask, LIME(Ribeiro, Singh, and Guestrin 2016), RISE(Petsiuk, Das, and Saenko 2018), Greydanus et al. (2018), and SARFA(Puri et al. 2020) for the PPO agent(Schulman et al. 2017; Nguyen 2020) trained in the Super Mario Bros.(Kauten 2018) environment. Saliency level ranges from blue(low) to red(high). Row 1: The chosen action is to move left. The explanation is that Mario is attempting to avoid the Piranha Plants on top of the green pipes. VisionMask accurately identified the two Piranha Plants on the pipes. While RISE and Greydanus highlighted almost the entire image and LIME and SARFA highlighted areas not directly related to the explanation. Row 2:The chosen action is to jump right with acceleration. The explanation is that Mario intends to land on top of the Koopas to clear his path and get reward. Again VisionMask clearly highlighted the two Koopas whereas the other saliency maps failed to identify the correct region.
Deep Reinforcement Learning (DRL) is a powerful technology in machine intelligence, widely used for many applications, including ChatGPT (Brown et al. 2020; Achiam et al. 2023) and Alpha Go Zero (Silver et al. 2017). It enables effective end-to-end training of decision models that map an agent’s observations to actions, aiming to maximizing the long term expected rewards. However, understanding a DRL agent’s decision-making process is challenging, due to the inherent lack of explainability in the high-dimensional, non-linear structure of its underlying Deep Neural Network (DNN). The lack of transparency undermines human users’ trust and acceptance, particularly in safety-critical scenarios, such as medical diagnosis and military operations. To ensure wide acceptance and efficient use of this technology, it is crucial to provide accurate and human-interpretable explanations. These considerations motivated the recent interests in Explainable Reinforcement Learning (XRL).
Some existing works integrate explanation generation into the agent model and train them together. We refer to these as model-based approaches. For example, Interpretable DQN (i-DQN) (Annasamy and Sycara 2019) implements a deep Q-network with an attention mechanism, using the attention weights between state and action embeddings to represent their importance. Saliency-guided Q-networks (SGQNs) (Bertoin et al. 2022) calculate the importance of inputs as a derivative of the Q-function and train the critic model to both minimize the value prediction error and maximize the fidelity of the explanation. Some other works moved away from DNN and adopt special policy architectures that are inherently more interpretable (Liu et al. 2018; Verma et al. 2018; Bastani, Pu, and Solar-Lezama 2018). However, these models usually experience performance loss compared to the DNN-based policy model due to the limitations imposed by the special structure. Those model-based approaches cannot provide explanations for a trained DRL agent, unless they retrain the agent using their specific models, which is prohibitively expensive. And the retrain process may compromise the agent’s performance and integrity as it strives to maintain a balance between the optimality and explainability.
One of the most widely used techniques to provide explanations for machine learning models is the attribution-based method, which assigns credits to input features based on their contribution to AI agent’s decisions. This technique was originally developed to explain image classification models in computer vision. It generates saliency maps by perturbing the input image to reveal their contributions to the classification process without modifying the classifier itself (Dabkowski and Gal 2017; Petsiuk, Das, and Saenko 2018; Ribeiro, Singh, and Guestrin 2016). Inspired by this, several works have explored to use attribution-based technique to explain DRL models while treating the agent as a black-box. Greydanus et al. (2018) perturb the input features using Gaussian blur at fixed intervals and measure how the perturbation changes the action probability. The difference in action probability before and after perturbation are used to generate a saliency-map for the importance of each input feature. Similarly, Iyer et al. (2018) and Puri et al. (2020) compute the importance score based on the differences in Q values before and after the perturbation. The latter also improve the accuracy of the saliency map by normalizing the effects of perturbation between target and non-target actions.
The perturbation-based methods do not alter the agent model, nor do they impose any specific requirements on it. Hence they are applicable to any trained DRL agents. However, these techniques have some significant limitations. First, to manage complexity, perturbations are performed according to a set of predefined rules, such as using Gaussian blur or occluding with a constant value, which severely limits the search space they can explore. For DRL agents that map visual input directly to actions, due to the large input dimensions, there is no guarantee that the predefined perturbation can find the set of pixels that trigger the agent decision. Secondly, perturbation-based methods do not learn from past experiences, nor do they improve performance by comparing similarities and differences across different input states and actions. Thirdly, each input feature (e.g., input pixel) is perturbed independently, with no guarantee that the joint state of the perturbed pixels represents meaningful objects. As shown in Table 1, the saliency maps generated by the perturbation-based method either consider all pixels equally important or identify regions in the image that do not contain meaning objects.
To address the aforementioned challenges, we propose VisionMask, a model trained end-to-end to generate attribution-based explanations for vision-based DRL without perturbation or altering the agent. VisionMask ouputs a saliency map that highlights the critical regions in the agent’s visual input responsible for triggering the agent’s action. The saliency map reveals the importance of each pixel in the decision making process, and will also be referred to as importance map or importance mask in the rest of the paper. VisionMask is trained using self-supervised contrastive learning. No human generated label is required. The training of VisionMask does not alter the agent model, thereby preserving the agent’s performance and integrity. The (fixed) policy model of the DRL agent is integrated into the training pipeline to generate the loss, and involved in the backpropagation to pass the gradient to the mask generation model. Therefore, each VisionMask model is trained to explain a specific DRL agent. The training process is agent agnostic, meaning that it works with any agent with a policy model that maps visual input to actions. VisionMasks’s learnable capability allows it to accumulate knowledge from previous experiences to improve future explanations. During testing time, VisionMask does not need to access agent’s internal state to generate the importance mask. The mask is generated by a forward pass of the VisionMask model.
As shown in Table 1, VisionMask accurately identifies the critical regions in the visual input that explains the agent’s action. Beyond providing precise explanations directly from the visual input, VisionMask also facilitates counterfactual analysis. By selectively activating or deactivating different critical regions, meaningful counterfactual scenarios can be generated and assessed. This is similar to the perturbation-based approach, however, at a much higher level.
The contribution of our work can be summarized as the following:
  • We present VisionMask, a novel attribution method for explainable visual reinforcement learning that generates action-specific importance masks. Generated using a single forward pass, the mask does not only highlight the regions in the visual input that explain the agent’s behavior but also facilitates counterfactual analysis.
  • VisionMask is model-agnostic in that it works with any vision-based DRL agent without modifying the agent model.
  • VisionMask is trained using self-supervised contrastive learning. No human intervention is needed for data labeling.
  • The fidelity of the explanation is measured by the agreement between the agent’s actions generated from the original visual input versus from the masked visual input. We quantify this agreement using accuracy, precision, recall and F1-score. In average, VisionMask outperforms other attribution-based explanation models by 30.08% on F1-score and 14.9% on insertion accuracy..
2 Background and related work
The nested structure and non-linear operation of DNN make it challenging for humans to understand how the outputs are derived from inputs. Some existing works address this challenge by explaining model outputs through input attributions (Dabkowski and Gal 2017; Petsiuk, Das, and Saenko 2018; Ribeiro, Singh, and Guestrin 2016; Fong and Vedaldi 2017). For example, Randomized Input Sampling for Explanation of Black-box Models (RISE) (Petsiuk, Das, and Saenko 2018) is a perturbation-based approach that explains neural image classifiers by applying randomly generated masks to the input image and assessing their impact. Local Interpretable Model-agnostic Explanations (LIME)(Ribeiro, Singh, and Guestrin 2016) tries to explain local instances by approximating them within a nearby vicinity using a linear decision model, where the explainable elements are superpixels (i.e., small image regions of similar pixels).
Built on top of the DNNs, the lack of transparency of the DRL agents (Samek, Wiegand, and Müller 2017; Ras et al. 2022; Angelov and Soares 2020) undermines humans trust and hinders their adoption. Several studies have addressed this challenging problems. According to (Qing et al. 2022), existing efforts can be categorized into four main approaches: model explaining, reward explaining, task explaining, and state explaining.
Model explaining relies on inherently interpretable model architectures or auxiliary reasoning mechanisms to generate explanations (Topin et al. 2021; Jiang and Luo 2019; Landajuela et al. 2021). However, these self-explanatory or symbolic models often suffer from decreased performance compared to state-of-the-art neural network based RL policies and may lack the representational power needed to learn more complex policies. Reward explanation typically involves using explicitly designed explainable reward functions to generate explanations (Juozapaitis et al. 2019; Ashwood, Jha, and Pillow 2022). Task explanation considers a policy as a composition of sub-task and action sequences, explaining the behavior in terms of relationships among sub-tasks (Shu, Xiong, and Socher 2017; Sodhani, Zhang, and Pineau 2021). These approaches often assume the existence of a reward decomposition scheme or a predefined sub-task partition mechanism, which may not hold true in all RL environments.
State explanations are based on agent observations from the environment. This approach determines the significance of explainable elements within a state observation in relation to action selection or reward determination. An explainable element in the state could be a small region in the visual input or semantic features in the environment state. The proposed VisionMask falls into this category.
State explaining methods can further be divided into three categories: attention-based, Shapley value-based and perturbation-based mechanisms, which are detailed below. Attention is a common approach used for explainable RL in several existing works (Peng, Riedl, and Ammanabrolu 2022; Xu et al. 2020; Annasamy and Sycara 2019). However, similar to i-DQN (Annasamy and Sycara 2019), these methods cannot explain a given DRL model because their attention model must be trained alongside the agent model. Shapley value (Shapley et al. 1953), a concept from game theory, has also been introduced into XRL. SVERL (Beechey, Smith, and Şimşek 2023) provides a theoretical analysis on how Shapley values can be applied to explain value and policy network in DRL. However, this approach is still in its early stages and assumes prior knowledge of the transition model of the environment, which is not realistic for realistic applications.
Perturbation-based methods (Greydanus et al. 2018; Iyer et al. 2018; Puri et al. 2020) compute saliency maps of the input features by comparing the action probability or value function before and after perturbation. Greydanus et al. (2018) perturbed the state with Gaussian blur at fixed intervals, while Iyer et al. (2018) identified objects and perturbed them with a constant gray value. Puri et al. (2020) improved on Greydanus et al. (2018) by removing the effect of perturbations from other irrelevant actions. As discussed in section 1, the performance of these techniques is constrained by predefined perturbation rules and the lack of knowledge accumulation. Furthermore, there is no guarantee that the perturbed input remains physically meaningful.
3 VisionMask
This section presents VisionMask, an innovative agent-agnostic framework for XRL. The primary goal is to produce action-wise importance-mask that attribute explainable elements in the current state to each possible action. The definition of explainable elements may vary across different applications. However, in this paper, we focus on identifying explanatory pixel regions in the agent’s visual input.
Figure 1: Architecture of VisionMask. In this example $a=1$ so $m_{1}$ is the active mask $m_{a}$.
3.1 Problem Formulation
We approach the task of VisionMask as an imitation learning problem based on Behavioral Cloning (BC) (Bain and Sammut 1995; Ross, Gordon, and Bagnell 2011). Formally, we define the environment as a Markov Decision Process (MDP) $\{\mathcal{S},\mathcal{A},P,R,\gamma\}$, where $\mathcal{S}$ represents the state space; $\mathcal{A}$ denotes the action space; the state transition function $P:\mathcal{S}\times\mathcal{A}\to\Delta(\mathcal{S})$ depicts the transition between states based on actions, where $\Delta(\mathcal{S})$ represents the set of probability distributions over states; the reward function $r:\mathcal{S}\times\mathcal{A}\to\mathbb{R}$ provides the immediate reward for state-action pairs; and $\gamma\in[0,1]$ and $\pi:\mathcal{S}\to\Delta(\mathcal{A})$ represent the discount factor and policy. Return $G$ is defined as $G=\sum_{k=0}^{\infty}\gamma^{k}R_{k+1}$, and the expected cumulative reward of a policy $\pi$ is $\mathbb{E}_{\pi}[G]=\mathbb{E}_{\pi}[\sum_{k=0}^{\infty}\gamma^{k}R_{k+1}]$, where the expectation is taken with respect to the initial state distribution, transition probabilities, and action probabilities determined by $\pi$.
VisionMask operates on a given pre-trained expert policy $\pi_{E}$ such that
$\pi_{E}\approx\pi^{*}=\operatorname*{arg\,max}_{\pi}\mathbb{E}_{\pi}[G],$
where $\pi^{*}$ is the optimal policy. We can obtain a dataset of expert demonstrations $\mathcal{D}_{E}=\{(s_{i},\pi_{E}(s_{i}))\}_{i=1}^{N}$ consisting of $N$ state-action pairs, from trajectories sampled while executing $\pi_{E}$ in the environment. Our goal is to learn an explainer $f_{\theta^{*}}$ that minimizes the loss
$\theta^{*}=\operatorname*{arg\,min}_{\theta}\sum_{(s,a)\in\mathcal{D}_{E}}L(a,s,\theta),$
where $L$ is the training loss function to be discussed in section 3.3. The explainer function $f_{\theta}:\mathcal{S}\to[0,1]^{K\times d_{s}}$ (with $d_{s}$ representing the state dimension and $K$ the number of candidate actions) provides the attributions of each action to each feature in the state $s$. In the case of visual DRL, $s$ is the visual input of the agent and $d_{s}$ is the number of pixels in $s$, i.e., $d_{s}=W\times{H}$, with $W$ and $H$ representing the width and height of the visual input. The value of each element in the attribution is bounded within the range $[0,1]$, with the $(i,j)$th element indicates the $j$th pixel’s attribution to the $i$th action. This attribution will then be overlaid onto the visual input, masking unimportant pixels while retaining important ones. The overlaid visual input will be analyzed by the policy model again. We will update parameters $\theta$ of explainer $f$ so that the overlaid visual input always produces the same action as the original one. The expert policy $\pi_{E}$ remains frozen during the training process of $f_{\theta}$, meaning its parameters are fixed.
3.2 Architecture
As shown in Fig. 1, the expert dataset $\mathcal{D}_{E}$ is collected using the expert policy $\pi_{E}$. From this dataset $\mathcal{D}_{E}$, state-action pairs $(s_{i},a_{i})$ are sampled. Subsequently, the explainer $f_{\theta}$ takes the state $s_{i}$ as input and produces a set of action-wise masks $M=\{m_{0},m_{1},....,m_{K-1}\}$, where $K$ denotes the number of candidate actions and $m_{i}\in{[0,1]^{W\times{H}}}$. A mask value $m_{i}[x,y]$ near 1 indicates that the pixel $(x,y)$ has significant contributions to the corresponding action $i$, whereas a value near 0 denotes irrelavance. This process shares similarities with image segmentation, where each pixel is attributed to an object or background. Hence, we structure the explainer $f_{\theta}$ akin to the widely used image segmentation model, DeepLabV3 (Chen et al. 2017), but retrain it with the expert demonstrations $\pi_{E}$. To make sure that the output masks are bounded to the range $[0,1]$, a sigmoid function is applied at the output of $f_{\theta}$
Following (Stalder et al. 2022), we partition the masks in $M$ into two sets: the active set and the inactive set. Unlike (Stalder et al. 2022), our active set has only one mask, $m_{a}$ where $a$ is the action selected by expert policy $\pi_{E}$ll the other masks, $m_{i}$, where $i\neq{a}$, belong to the inactive set. Each mask $m_{i}$ contains the attributions for the current agent state that explain the corresponding action $i$. The explainer will be trained to produce masks for pixels that contribute the most to the policy’s action selection.
We aggregate the masks in the inactive set into a non-target mask $n$ by taking the maximum value pixel-wise. We also calculate a complement mask $\tilde{m_{a}}=\mathbf{1}-m_{a}$ representing the less important regions for the selected action $a$. Then the masks $m_{a}$ and $\tilde{m_{a}}$ are overlaid onto the original state $s$ to generate two masked states $s_{m}$ and $s_{\tilde{m}}$ using the following overlay function:
$s_{m}=s\odot m_{a}+r\odot(\tilde{m_{a}}),s_{\tilde{m}}=s\odot\tilde{m_{a}}+r\odot(m_{a})$
where $\odot$ is Hadamard Product and $r$ is a reference value. The final step is to feed these masked states $s_{m}$ and $s_{\tilde{m}}$ through the expert policy $\pi_{E}$ to generates action distributions $\mathbf{p}$ and $\mathbf{\tilde{p}}$, respectively:
$\mathbf{p}=\pi_{E}(\mathbf{s}_{\mathbf{m}}),\mathbf{\tilde{p}}=\pi_{E}(\mathbf{s}_{\mathbf{\tilde{m}}}).$
The model $f_{\theta}$ is optimized so that $\mathbf{p}$ matches the action distribution chosen for the unmasked state $s$, while $\mathbf{\tilde{p}}$ does not.
It is worth mentioning that the agent may be sensitive to the reference value in the overlay function. Numerous options exist for the reference value, such as setting the pixel to zero, assigning a constant value, blurring the pixel, or cropping it. In this study, we opt for a constant value as the reference value, based on the sensitivity of the agent to zero and blurring, as indicated by experimental results. We use different reference values for different environments (See Appendix 4).
3.3 Training Loss
To train the explainer to capture the most important regions in the state, we carefully designed the training loss function $\mathcal{L}$ as follows:
$\mathcal{L}(\mathbf{s},\mathbf{a},\theta,\mathbf{n})$ $=\mathcal{L}_{bc}(\mathbf{s},\mathbf{a})+\lambda_{e}\mathcal{L}_{e}(\mathbf{s},\mathbf{a})+\lambda_{avg}\mathcal{L}_{avg}(\mathbf{s},\mathbf{a},\mathbf{n})$
$+\lambda_{\text{smooth}}\mathcal{L}_{\text{smooth}}(\mathbf{s},\mathbf{a},\mathbf{n})+\lambda_{L2}\lVert\theta\rVert^{2}_{2},$
where $\mathcal{L}_{bc}(\mathbf{s},\mathbf{a})$ is Behavior Cloning loss, $\mathcal{L}_{e}(\mathbf{s},\mathbf{a})$ is entropy loss, $\mathcal{L}_{avg}(\mathbf{s},\mathbf{a},\mathbf{n})$ is the mask average loss, $\mathcal{L}_{\text{smooth}}(\mathbf{s},\mathbf{a},\mathbf{n})$ is the smoothness loss, $\lVert\theta\rVert^{2}_{2}$ is $L_{2}$ regularization and $\lambda_{e},\lambda_{avg},\lambda_{\text{smooth}},\lambda_{L2}$ are regularization hyper-parameters.
Behavioral Cloning loss $\mathcal{L}_{bc}$. Our goal is to train the explainer to select regions in the visual input that that are sufficient and necessary for the agent to choose action $\mathbf{a}$. Hence, we design the behavior cloning loss as cross entropy loss between the action distribution $\mathbf{p}$ predicted for the overlaid state $s_{m}$ and the target $\mathbf{a}$ chosen for the original state $s$:
$\mathcal{L}_{bc}(\mathbf{s},\mathbf{a})=-\frac{1}{K}\sum_{k=1}^{K}\llbracket k=\mathbf{a}\rrbracket\log(\mathbf{p}[k]).$
Here $\llbracket k=\mathbf{a}\rrbracket$ denotes the indicator function which returns 1 if $k$ is the same as label $\mathbf{a}$, and 0 otherwise.
Negative Entropy loss $\mathcal{L}_{e}$. To ensure that the visual input regions selected by $m_{a}$ is sufficient and necessary for the agent to make decisions, we also need to make sure that the unselected region, i.e., $s_{\tilde{m_{a}}}$, does not provide useful information for action selection, hence the action distribution $\mathbf{\tilde{p}}$ should follow a uniform distribution. Motivated by this rationale, we define negative entropy loss regarding $\tilde{\mathbf{m}}$ as the following:
$\mathcal{L}_{e}(\mathbf{s},\mathbf{a})=\frac{1}{K}\sum^{K}_{k=1}\tilde{\mathbf{p}}[k]\log\tilde{\mathbf{p}}[k].$
Mask average loss $\mathcal{L}_{avg}$. A low effort way to minimize $\mathcal{L}_{bc}$ and $\mathcal{L}_{e}$ is to include all pixels in the importance mask, $m_{a}$, and no pixel in the complement mask, $\tilde{m_{a}}$, which obviously is not a valid solution. We need to ensure that each importance mask only consists a small number of crucial pixels. Thus, we define $\mathcal{L}_{avg}$ as follows:
$\mathcal{L}_{avg}(\mathbf{s},\mathbf{a},\mathbf{n})=\frac{1}{W\times H}\sum_{i,j}\mathbf{m}[i,j]+\frac{1}{W\times H}\sum_{i,j}\mathbf{n}[i,j],$
where, $\mathbf{n}$ represents the non-target mask aggregated from the inactive set. By minimizing the mask average loss, we limit the number of pixels selected by the masks.
Smoothness loss $\mathcal{L}_{\text{smooth}}$. Finally, we minimize the Variation between adjacent pixels to generate smoother masks:
$\mathcal{TV}(x)$ $=\frac{1}{W\times H}\sum_{i,j}(|x[i,j]-x[i+1,j]|$
$+|x[i,j]-x[i,j+1]|)$
$\mathcal{L}_{\text{smooth}}(\mathbf{s},\mathbf{a},\mathbf{n})=\mathcal{TV}(m)+\mathcal{TV}(n)$
The overall architecture of VisionMask is given in Fig. 1. The loss will backpropagate through the frozen expert model $\pi_{E}$ to train the explainer model $f_{\theta}$.
4 Experiments
Env Metric VisionMask LIME RISE Greydanus SARFA
Mario Acc. $\mathbf{0.9076}\pm\scriptstyle{0.0124}$ $0.7868\pm\scriptstyle{0.0027}$ $0.8053\pm\scriptstyle{0.0118}$ 0.1692 0.1692
Precision $\mathbf{0.7109}\pm\scriptstyle{0.0189}$ $0.2448\pm\scriptstyle{0.0200}$ $0.3364\pm\scriptstyle{0.0046}$ 0.6154 0.6154
Recall $\mathbf{0.7005}\pm\scriptstyle{0.0020}$ $0.2467\pm\scriptstyle{0.0042}$ $0.3245\pm\scriptstyle{0.0069}$ 0.1752 0.1800
F1-score $\mathbf{0.6995}\pm\scriptstyle{0.0058}$ $0.2277\pm\scriptstyle{0.0056}$ $0.3046\pm\scriptstyle{0.0082}$ 0.2764 0.2753
Enduro Acc. $\mathbf{0.9061}\pm\scriptstyle{0.0077}$ $0.8811\pm\scriptstyle{0.0087}$ $0.9033\pm\scriptstyle{0.0047}$ 0.2744 0.2790
Precision $0.6058\pm\scriptstyle{0.0084}$ $0.4167\pm\scriptstyle{0.0062}$ $0.5363\pm\scriptstyle{0.0054}$ $\mathbf{0.6449}$ 0.6383
Recall $\mathbf{0.5535}\pm\scriptstyle{0.0049}$ $0.4551\pm\scriptstyle{0.0046}$ $0.5245\pm\scriptstyle{0.0072}$ 0.0839 0.0889
F1-score $\mathbf{0.5060}\pm\scriptstyle{0.0082}$ $0.3809\pm\scriptstyle{0.0018}$ $0.4738\pm\scriptstyle{0.0032}$ 0.1283 0.1418
Seaquest Acc. $\mathbf{0.9531}\pm\scriptstyle{0.0030}$ $0.9094\pm\scriptstyle{0.0079}$ $0.9201\pm\scriptstyle{0.0051}$ 0.4463 0.4476
Precision $0.5809\pm\scriptstyle{0.0022}$ $0.2647\pm\scriptstyle{0.0066}$ $0.4631\pm\scriptstyle{0.004}$ 0.8083 $\mathbf{0.8371}$
Recall $\mathbf{0.4880}\pm\scriptstyle{0.0013}$ $0.2423\pm\scriptstyle{0.0082}$ $0.2830\pm\scriptstyle{0.0047}$ 0.0945 0.0983
F1-score $\mathbf{0.4714}\pm\scriptstyle{0.0050}$ $0.2073\pm\scriptstyle{0.01443}$ $0.2733\pm\scriptstyle{0.0110}$ 0.1640 0.1701
MsPacman Acc. $\mathbf{0.9768}\pm\scriptstyle{0.0177}$ $0.8451\pm\scriptstyle{0.0294}$ $0.9264\pm\scriptstyle{0.0220}$ 0.4616 0.4235
Precision $0.7671\pm\scriptstyle{0.0160}$ $0.1878\pm\scriptstyle{0.0298}$ $0.5590\pm\scriptstyle{0.0097}$ 0.8255 $\mathbf{0.9102}$
Recall $\mathbf{0.8658}\pm\scriptstyle{0.0039}$ $0.1726\pm\scriptstyle{0.0036}$ $0.5587\pm\scriptstyle{0.0035}$ 0.2189 0.1913
F1-score $\mathbf{0.7699}\pm\scriptstyle{0.0096}$ $0.1597\pm\scriptstyle{0.0086}$ $0.5122\pm\scriptstyle{0.0078}$ 0.3307 0.3067
Table 2: Fidelity scores on four environments and confusion matrices in the Mario environment. We run each experiment 3 times with different random seeds and report the mean and standard deviation.
We evaluated VisionMask across four different environments and compared its performance against four baseline methods.
Environment Selection. Four agent-centered RL environments were selected to showcase the effectiveness of VisionMask: Super Mario Bros (SMB)(Kauten 2018), and the Atari games Seaquest, Enduro, and MsPacman. We select these environments following existing works (Puri et al. 2020; Greydanus et al. 2018).
Baseline Selection. We compared VisionMask’s performance against four perturbation-based explainer, LIME (Ribeiro, Singh, and Guestrin 2016) RISE (Petsiuk, Das, and Saenko 2018), Greydanus et al. (2018) and SARFA(Puri et al. 2020). For LIME, we used the original public implementation with the default configuration available from (Ribeiro 2017). For RISE, we adopt the public implementation from torchray (Fong, Patrick, and Vedaldi 2019). For Greydanus et al. (2018) and SARFA, we used the respective public implementation provided by the authors. To ensure a fair comparison, we did not include any explainers that require retraining the agent model.
Expert Policy. For reproducibility, we did not train the expert policies ourselves; instead, we used publicly available implementations. In the SMB environment, We apply the trained PPO (Schulman et al. 2017) policy from (Nguyen 2020). In the Atari environments, we used three trained PPO policies from Stable-Baselines3 (Raffin et al. 2021) .
4.1 Assessment of Explanation Fidelity
We quantify fidelity using four metrics: accuracy, precision, recall, and F1-score. Fidelity measures the agreement between the agent’s action selected for the original visual input and the action selected for the masked visual input. In this context, we define the accuracy as the following:
$\text{Accuracy}=\frac{1}{|\mathcal{D}_{\text{test}}|}\sum_{i=1}^{|\mathcal{D}_{\text{test}}|}\llbracket\pi_{E}(s_{i})=\pi_{E}(s_{i}^{\prime})\rrbracket,$
[…]
Values of these fidelity metrics are shown in Table 2. VisionMask consistently scores higher than LIME and RISE. Greydanus and SARFA have a higher precision than VisionMask, however, a much lower recall and F1-score. This is probably because they included almost all pixels from the visual input in the importance map. They also have a lower accuracy compared to VisionMask.
Refer to caption Refer to caption
Original state VisionMask
Refer to caption Refer to caption
Insertion curve Deletion curve
Figure 2: Examples of insertion and deletion AUC curve
Env Method 0.25 0.5 1.0
$\downarrow$ del $\uparrow$ ins $\downarrow$ del $\uparrow$ ins $\downarrow$ del $\uparrow$ ins
Mario VisionMask 0.2044 ±0.1891 0.4156 ±0.2937 0.1813 ±0.1730 0.4142 ±0.2788 0.1714 ±0.1577 0.4216 ±0.2622
LIME 0.3447 ±0.2664 0.2314 ±0.2189 0.3231 ±0.2568 0.2446 ±0.2232 0.2693 ±0.1958 0.3024 ±0.2107
RISE 0.2750 ±0.2510 0.2940 ±0.2624 0.2331 ±0.2326 0.3317 ±0.2761 0.1892 ±0.1897 0.3796 ±0.2678
Greydanus 0.5708 ±0.2109 0.1963 ±0.1658 0.5671 ±0.2146 0.1952 ±0.1650 0.5634 ±0.2169 0.2018 ±0.1691
SARFA 0.5579 ±0.2223 0.2071 ±0.1756 0.5565 ±0.2235 0.2117 ±0.18 0.5556 ±0.2246 0.2143 ±0.1866
Enduro VisionMask 0.3693 ±0.2854 0.3952 ±0.3004 0.3589 ±0.2911 0.3979 ±0.2995 0.3491 ±0.2948 0.4021 ±0.2941
LIME 0.3780 ±0.2886 0.3636 ±0.3025 0.3723 ±0.2914 0.3680 ±0.3002 0.3629 ±0.2937 0.3799 ±0.2883
RISE 0.4081 ±0.2818 0.3585 ±0.3066 0.3970 ±0.2861 0.3714 ±0.2938 0.3759 ±0.2912 0.3906 ±0.2785
Greydanus 0.3762 ±0.2988 0.3312 ±0.3194 0.3626 ±0.3018 0.3317 ±0.3192 0.3457 ±0.3075 0.3329 ±0.3188
SARFA 0.3472 ±0.3060 0.3327 ±0.3188 0.3378 ±0.3114 0.3336 ±0.3186 0.3331 ±0.3157 0.3340 ±0.3186
Seaquest VisionMask 0.0721 ±0.0291 0.3256 ±0.2596 0.0684 ±0.0289 0.3331 ±0.2650 0.0652 ±0.0268 0.3430 ±0.2635
LIME 0.1982 ±0.1871 0.1397 ±0.1662 0.1909 ±0.1865 0.1425 ±0.1672 0.1434 ±0.1214 0.2141 ±0.1812
RISE 0.1198 ±0.0797 0.2422 ±0.2560 0.0987 ±0.0658 0.2638 ±0.2654 0.0809 ±0.0505 0.2896 ±0.2660
Greydanus 0.2939 ±0.2381 0.0658 ±0.0345 0.2165 ±0.1834 0.0708 ±0.0542 0.1439 ±0.0980 0.0788 ±0.0825
SARFA 0.1133 ±0.0732 0.0819 ±0.0978 0.0888 ±0.0418 0.0851 ±0.1104 0.0764 ±0.0307 0.0869 ±0.1177
MsPacman VisionMask 0.2251 ±0.1051 0.6350 ±0.3056 0.2037 ±0.1255 0.6231 ±0.3337 0.1956 ±0.1217 0.6278 ±0.2778
LIME 0.4831 ±0.2700 0.1901 ±0.1708 0.4348 ±0.2455 0.2346 ±0.1875 0.3265 ±0.1801 0.3771 ±0.1890
RISE 0.4270 ±0.3140 0.2299 ±0.1220 0.3512 ±0.2995 0.2937 ±0.1325 0.2621 ±0.2273 0.4577 ±0.1394
Greydanus 0.5669 ±0.2476 0.1285 ±0.0646 0.4903 ±0.2518 0.1336 ±0.0843 0.3905 ±0.2281 0.1417 ±0.1029
SARFA 0.3511 ±0.2528 0.1463 ±0.1168 0.2524 ±0.1573 0.1531 ±0.1328 0.1880 ±0.0873 0.1561 ±0.1425
Table 3: Insertion and deletion metrics for fractions 0.25, 0.5, and 1.0. The symbol $\uparrow$ (resp. $\downarrow$) indicates that higher (resp. lower) values are better. We report the mean $\pm$ standard deviation over testing samples.
4.2 Assessment of Effectiveness
An effective explainer should precisely identify the exact input regions that contribute to the expert decision, without including irrelevant areas. To quantify the effectiveness of explanations, We adopt the insertion and deletion metrics from RISE (Petsiuk, Das, and Saenko 2018). The rationale behind these metrics is as follows: If certain pixels in the visual state are crucial to the action (i.e., the true cause of the action), deleting them from the state or inserting them into the background should result in a corresponding decrease or increase in the probability of the specific action. By plotting the relationship between the probability of the target action and the number of pixels added or removed, we obtain an insertion curve and a deletion curve. The Area Under the Curve (AUC) of those plots represents the effectiveness of the explanation. A higher AUC for the insertion curve and a lower AUC for the deletion curve stand for more effective explanation. The order of insertion and deletion affects the result. We sort the pixels based on their values in the importance map and insert or delete them in descending order of their importance. Figure 2 shows an example of insertion and deletion curves with corresponding AUC values displayed at the top right corner. See Appendix D for the pseudocode of each metric.
It is worth noting that the effectiveness varies as different number of pixels are inserted or deleted from the original input. Instead of showing the actual curve for each explainer under different games, we report the AUC under 25%, 50%, and 100% level of insertion and deletion. The results are given in Table 3. VisionMask outperforms all other methods across all environments, with an average insertion score that is 14.9% higher and a deletion score that is 8.8% lower compared to the other methods.
4.3 Counterfactual Explanations
Original state VisionMask Counterfactual
[Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
Move left Jump right
[Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
Left Left + acceleration
[Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
Down Fire Down right
Figure 3: Counterfactual analysis with VisionMask.
By removing the most important regions identified by VisionMask from the visual input, we can create scenarios for counterfacutal analysis. Table 3 provides three examples of this approach. In the first example, VisionMask correctly identified the Piranha Plants as the reason the agent chose to move left. After removing the Piranha Plant by replacing the highlighted pixels with reference value and run the modified image by the agent again, we observed that the agent’s action changed to jumping right, as the threat from the Piranha Plant no longer existed. In the second example, VisionMask correctly identified the edge of the road and a car in front as the reason why the agent chose to move left (with current speed). After we removed the car in front the agent chose to move left with acceleration. In the third example, VisionMask identified the shark in the bottom left as the cause of the downward and fire action. When the corresponding shark was removed, the agentno longer chose to fire. Appendix I, J and K contain more examples. We believe such explanations can further help users to understand the RL agent’s behavior.
5 Conclusion
We presented VisionMask, an agent-agnostic DRL explanation model trained in self-supervised contrastive learning. VisionMask generates explanations with higher fidelity and better effectiveness compared to existing attribute-based methods. It is our future plan to extend this approach to multi-modality input and couple the visual explanation generated by the VisionMake with other information such as agent’s long-term goals and future rewards.
Acknowledgements
This research is partially supported by the Air Force Office of Scientific Research (AFOSR), under contract FA9550-24-1-0078. The paper was received and approved for public release by AFRL on May 28th 2024, case number AFRL-2024-2908 . Any Opinions, findings, and conclusions or recommendations expressed in this material are those of the authors and do not necessarily reflect the views of AFRLor its contractors.
Appendix A Technical Appendix
Appendix B Meanings of actions across four environments
In Table 4, we present the list of actions and their corresponding meanings for the four environments evaluated in the experiment section.
Value Meaning Value Meaning
0 Stay in place 0 No operation
1 Move right 1 Accelerate
2 Jump right 2 Move right
3 Move right with acceleration 3 Move left
4 Jump right with acceleration 4 Decelerate
5 Jump 5 Move right and decelerate
6 Move left 6 Move left and decelerate
7 Accelerate and move right
8 Accelerate and move left
(a) Mario environment (b) Atari Enduro environment
Value Meaning Value Meaning
0 No operation 0 No operation
1 Fire 1 Move up
2 Move up 2 Move right
3 Move right 3 Move left
4 Move left 4 Move down
5 Move down 5 Move up-right
6 Move up-right 6 Move up-left
7 Move up-left 7 Move down-right
8 Move down-right 8 Move down-left
9 Move down-left
10 Fire and move up
11 Fire and move right
12 Fire and move left
13 Fire and move down
14 Fire and move up-right
15 Fire and move up-left
16 Fire and move down-right
17 Fire and move down-left
(c) Atari Seaquest environment (d) Atari MsPacman environment
Table 4: Action meanings for SMB (Nguyen 2020), Atari Enduro, Seaquest and MsPacman (Brockman et al. 2016)
Appendix C Reference values used across different environments
In our experiment, we found that the PPO agents we adopted from Stable Baselines3 (Raffin et al. 2021) and Mario-PPO (Nguyen 2020) are sensitive to reference values. We tried using a constant value (zero) and blur as reference values but did not obtain clear and reasonable importance masks. However, using reference values that are visually similar to the environment background gave us better results. In Figure 4, we list the reference values we used in each environment. Specifically, for the SMB and Atari Enduro environments, we use the background color of the content in each state as the reference value. Note that the reference value for different states may vary, as the background content of each state can be different. For the Atari Seaquest environment, since all the states share the same background, we use a single reference value for each state, which combines the empty sea background and the oxygen bar background. For the Ms. Pacman environment, we simply used a constant value of the background color blue as the reference value.
Figure 4: Reference values for four environment evaluated. The content column represents a sub-figure of the original state that contains the objects and background. The last column displays the corresponding reference value for each environment.
Env Original state Content Reference value
SMB [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
Enduro [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
Seaquest [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
MsPacman [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
Appendix D Algorithms for insertion and deletion metrics
Algorithms 1-2 present the pseudocode for evaluating the insertion and deletion metrics of (Petsiuk, Das, and Saenko 2018).
Algorithm 1 Evaluate insertion metric
1:Input: target mask $m$, state $s$, action $a$, expert policy $\pi_{E}$, fraction of pixels inserted $\alpha$, Number of pixels $N$
2:Initialize $m^{\prime}\leftarrow\text{zeros}(m)$, $p\leftarrow\emptyset$
3:$\mathcal{I}\leftarrow\text{indices}(\text{sort}(\text{flatten}(m)))$
4:for $k\leftarrow 0$ to $\lfloor\alpha N\rfloor-1$ do
5:     $(x,y)\leftarrow\mathcal{I}[k]$, $m^{\prime}[x,y]\leftarrow 1.0$
6:     $p\leftarrow p\cup\pi_{E}(s\odot m^{\prime}+r\odot(1-m^{\prime}))$
7:end for
8:return $\text{AUC}(p)$
Algorithm 2 Evaluate deletion metric
1:Input: target mask $m$, state $s$, action $a$, expert policy $\pi_{E}$, fraction of pixels deleted $\alpha$, Number of pixels $N$
2:Initialize $m^{\prime}\leftarrow\text{ones}(m)$, $p\leftarrow\emptyset$
3:$\mathcal{I}\leftarrow\text{indices}(\text{sort}(\text{flatten}(m)))$
4:for $k\leftarrow 0$ to $\lfloor\alpha N\rfloor-1$ do
5:     $(x,y)\leftarrow\mathcal{I}[k]$, $m^{\prime}[x,y]\leftarrow 0.0$
6:     $p\leftarrow p\cup\pi_{E}(s\odot m^{\prime}+r\odot(1-m^{\prime}))$
7:end for
8:return $\text{AUC}(p)$
Appendix E Object importance percentage calculation
[…]
Appendix F Time efficiency
To assess the time efficiency of mask generation, we compare the total time taken for mask generation using VisionMask against that of two baseline methods on testing datasets from four environments. Note that the time we measure does not include the training time for VisionMask. We can see that LIME and RISE are slower than VisionMask. All the experiments are conducted on A40 GPU.
Env VisionMask LIME RISE
SMB 00:00:10 25:45:02 00:51:41
Enduro 00:00:10 1:05:19 00:08:14
Seaquest 00:00:12 1:34:49 00:10:58
MsPacman 00:00:10 3:02:28 00:08:59
Table 5: Time taken to generate importance masks for three methods across four environments. The time format is hh:mm:ss. The testing dataset for SMB contains 970 states, while Atari Enduro, Seaquest, and Ms. Pacman contain 4000 states each.
Appendix G Experimental details
Dataset details
Recall from Section 3.1 that we collected the expert dataset $\pi_{E}$ with the agent being explained as expert in four envrionments. In the SMB environment(Kauten 2018), $\pi_{E}^{\text{Mario}}$ comprises a total of 9,692 state-action pairs. In Atari Enduro, Seaquest and MsPacman (Brockman et al. 2016) environment, $\pi_{E}^{\text{Atari}}$ includes 10,000 total state-action pairs. We utilize a 80%-10%-10% train-valid-test split on all expert dataset $\pi_{E}$.
Pre-processing data
We transform the RGB input to greyscale with the following function then resize it to $84\times 84$:
1def rgb_to_grayscale(img, num_output_channels: int = 1):
2 if img.ndim < 3:
3 raise TypeError(f"Input image tensor should have at least 3 dimensions, but found {img.ndim}")
4 # _assert_channels(img, [1, 3])
5
6 if num_output_channels not in (1, 3):
7 raise ValueError("num_output_channels should be either 1 or 3")
8
9 if img.shape[-3] == 3:
10 r, g, b = img.unbind(dim=-3)
11 # This implementation closely follows the TF one:
12 # https://github.com/tensorflow/tensorflow/blob/v2.3.0/tensorflow/python/ops/image_ops_impl.py#L2105-L2138
13 l_img = (0.299 * r + 0.587 * g + 0.114 * b).to(img.dtype)
14 l_img = l_img.unsqueeze(dim=-3)
15 else:
16 l_img = img.clone()
17
18 if num_output_channels == 3:
19 return l_img.expand(img.shape)
20
21 return l_img
Model details
We utilize the segmentation model DeeplabV3 as the $f_{\theta}$ model in Section 3. See the Github repo of this model from  [^1] https://github.com/pytorch/pytorch. As mentioned in the Experiment section 4, we adopt the pretrained PPO model from Stable baselines 3 (Raffin et al. 2021) as the agent to be explained. Check the Hugging Face repositories for detailed information on the PPO agents for Enduro  [^2] https://huggingface.co/sb3/ppo-EnduroNoFrameskip-v4, Seaquest  [^3] https://huggingface.co/sb3/ppo-SeaquestNoFrameskip-v4, and MsPacman  [^4] https://huggingface.co/sb3/ppo-MsPacmanNoFrameskip-v4.
Training details
We conduct the training process on a server with an A40 GPU, using 50 epochs for SMB, 36 epochs for Enduro, 52 epochs for Seaquest, and 136 epochs for Ms. Pac-Man. The optimization is done using Adam. See Table 6 for hyperparameters.
Hyperparameter value
Learning rate 0.00001
Training batch size 16
Testing batch size 1
Validating batch size 16
$\lambda_{avg}$ 0.3
$\lambda_{e}$ 1.0
$\lambda_{\text{smooth}}$ 1.0
$\lambda_{L2}$ 0.01
Table 6: Hyperparameters
Three random seeds used for training: 42, 13, 62.
Appendix H Code
Code can be found at: https://anonymous.4open.science/r/I6eV3cksE5AgZ7-EB88
Appendix I Counterfactual explanations of Super Mario Bros environment
In this section, we show more counterfactual explanations on the Super Mario Bros(Kauten 2018) environment. In Table 7 below, each row represents a random state from the environment. The original state column shows the state, the VisionMask column displays the importance map of the target action (label), and the Counterfactuals column presents the corresponding counterfactuals constructed by removing the most important regions indicated by the importance map. We detail the counterfactual explanations in Table 7 below:
  • In the first row, Mario encounters the monster Goomba and two pipes, with the bush and Goomba receiving an importance score of 31.62%, while the right pipe receives 29.2%. Counterfactual scenarios involving the removal of the Goomba and the top section of the second green pipe both result in a change of action from accelerating while jumping right to simply jumping, which may indicate that those regions are the cause of action 4 (jump right with acceleration).
  • In the second row, Mario is positioned between two brown bricks, with the left brick receiving an importance score of 25.23%, the right brick receiving 29.15%, and Mario and the bush each receiving 25.23%. It’s worth noting that jumping is necessary in the original state for Mario to pass these obstacles, but it becomes unnecessary in the counterfactual scenario. This observation suggests that the bricks may be the cause of Action 4, as when they are removed, jumping is no longer necessary.
  • The third row is identical to the example presented in the counterfactual explanation section of the Super Mario Bros (SMB) environment in the paper.
  • In the fourth column, Mario encounters a brick wall that only requires a jump (without acceleration) to pass through, with the brick wall receiving an importance score of 28.78%. The first counterfactual, removing the brick, indicates a change in action from Action 2 (right jump) to Action 4. Knowing that a right jump with acceleration (Action 4) is necessary to safely jump over the gap and land on the next platform, one can easily identify the brick wall as the cause of the action. In addition, the observation that the second counterfactual of removing the ground behind Mario did not change the action, which in turn supports our previous explanation.
Note that all the object importance percentages are calculated following E.
Table 7: Counterfactual explanations of SMB environment.
Original state VisionMask Counterfactuals
[Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
Action 4 Action 5 Action 5
Jump right with acceleration Jump Jump
[Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
Action 4 Action 3
Jump right with acceleration Move right with acceleration
[Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
Action 6 Action 4 Action 1
Move left Jump right with acceleration Move right
[Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
Action 2 Action 4 Action 2
Jump right Jump right with acceleration Jump right
Appendix J Counterfactual explanations of Atari Enduro environment
In this section, we presents counterfactual explanations on the Atari Enduro environment. In Table 8 below, each row represents a random state from the environment. The original state column shows the state, the VisionMask column displays the importance map of the target action (label), and the Counterfactuals column presents the corresponding counterfactuals constructed by removing the most important regions indicated by the importance map. We detailed the counterfactual explanations in Table 8 below:
  • In the first row, the agent attempts to move left to avoid a collision with the incoming blue car, which receives an importance score of 8.33% as indicated by the importance map of second column. The counterfactual scenario involving the removal of the blue car leads to a change in action from moving left to accelerating while moving left. It’s important to note that accelerating while moving left could still result in a collision with the blue car in the original state.
  • In the second row, the agent attempts to overtake the blue car on the right, which receives an importance score of 11.25%. The corresponding counterfactual scenario of removing the blue car indicates a transition of action from accelerating to moving right and accelerating. Given that moving right with accelerating will cause a collision with that car, we may attribute the accelerating action to the blue car.
  • In the third row, a collision occurs in the original state with Action 3 (moving left). The agent tries to move left to avoid the collision but fails, where the collided blue car receives an importance score of 16.10%. By removing the blue (collided) car from the original state, the action in that counterfactual scenario changes to moving left with accelerating One may explain the collided blue car as the main cause of the action (moving left) in this counterfactual scenario.
  • In the last row, the agent tries to turn right, where the race track receives an importance score of 27.58%. In the counterfacutal scenario, the action changes from moving right with acceleration to moving left with acceleration when most of the race track is removed from original state. It’s obvious that the right directioned race track is the main cause of action of truing right with accelerating.
Table 8: Counterfactual explanations of Atari Enduro environment. Note that the collision happens in the original state in the third row.
Original state VisionMask Counterfactuals
[Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
Action 3: left Action 8: left + accelerate
[Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
Action 1: accelerate Action 7: right + accelerate
[Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
Action 3: left Action 8: left + accelerate
[Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
Action 7: right + accelerate Action 8: left + accelerate
Appendix K Counterfactual explanations of Atari Seaquest environment
In this section, we presents counterfactual explanations on the Atari Seaquest environment. In Table 9 below, each row represents a random state from the environment. The original state column shows the state, the VisionMask column displays the importance map of the target action (label), and the Counterfactuals column presents the corresponding counterfactuals constructed by removing the most important regions indicated by the importance map. We detailed the counterfactual explanations in Table 9 below:
  • In the first row, the agent attempts to attack the shark in the bottom left corner, marked in green, while the shark receives an importance score of 5.65%. In the counterfactual scenario where the shark is removed from the original state, the resulting action shifts from moving down and firing to moving down and to the right. The absence of the shark reasonably leads to a shift in action from firing to moving right, indicating that the presence of the shark likely prompted the original action.
  • The second row is similar to first row where the counterfactual of removing shark could incur the action shifting from fire to a movement.
  • In the last row, the agent’s submarine is about to start diving with a full tank of oxygen, where the oxygen bar receives most of the importance score, 53.51%. Counterfactual of removing the half of the oxygen bar making the action change from moving down, left and fire to moving left to stay on the surface. The oxygen level can easily be explained as the main cause in that counterfactual scenario.
Table 9: Counterfactual explanations of Atari Seaquest environment
Original state VisionMask Counterfactuals
[Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
Action 13: down + fire Action 8: down + right
[Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
Action 10: up + fire Action 6: up + right
[Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
Action 17: down + left + fire Action 4: left

### v2

@@
-Why the Agent Made that Decision: Explaining Deep Reinforcement Learning with Vision Masks
+Why the Agent Made that Decision: Contrastive Explanation Learning for Reinforcement Learning
 Abstract
-Due to the inherent lack of transparency in deep neural networks, it is challenging for deep reinforcement learning (DRL) agents to gain trust and acceptance from users, especially in safety-critical applications such as medical diagnosis and military operations.
-Existing methods for explaining an agent’s decision either require to retrain the agent using models that support explanation generation or rely on perturbation-based techniques to reveal the significance of different input features in the decision making process.
-However, retraining the agent may compromise its integrity and performance, while perturbation-based methods have limited performance and lack knowledge accumulation or learning capabilities.
-Moreover, since each perturbation is performed independently, the joint state of the perturbed inputs may not be physically meaningful.
-To address these challenges, we introduce VisionMask, a standalone explanation model trained end-to-end to identify the most critical regions in the agent’s visual input that can explain its actions.
-VisionMask is trained in a self-supervised manner without relying on human-generated labels.
-Importantly, its training does not alter the agent model, hence preserving the agent’s performance and integrity.
-We evaluate VisionMask on Super Mario Bros (SMB) and three Atari games.
-Compared to existing methods, VisionMask achieves a 14.9% higher insertion accuracy and a 30.08% higher F1-Score in reproducing original actions from the selected visual explanations.
-We also present examples illustrating how VisionMask can be used for counterfactual analysis.
+Reinforcement learning (RL) has demonstrated remarkable success in solving complex decision-making problems, yet its adoption in critical domains is hindered by the lack of interpretability in its decision-making processes.
+Existing explainable AI (xAI) approaches often fail to provide meaningful explanations for RL agents, particularly because they overlook the contrastive nature of human reasoning—answering ”why this action instead of that one?” To address this gap, we propose a novel framework of contrastive learning to explain RL selected actions, named VisionMask.
+VisionMask is trained to generate explanations by explicitly contrasting the agent’s chosen action with alternative actions in a given state using a self-supervised manner.
+We demonstrate the efficacy of our method through experiments across diverse RL environments, evaluating it in terms of faithfulness, robustness and complexity.
+Our results show that VisionMask significantly improves human understanding of agent behavior while maintaining accuracy and fidelity.
+Furthermore, we present examples illustrating how VisionMask can be used for counterfactual analysis.
+This work bridges the gap between RL and xAI, paving the way for safer and more interpretable RL systems.
 1 Introduction
-Original State LIME RISE Greydanus et al. SARFA VisionMask
-[Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
-[Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
-Table 1: Visual comparison of saliency maps generated by VisionMask, LIME(Ribeiro, Singh, and Guestrin 2016), RISE(Petsiuk, Das, and Saenko 2018), Greydanus et al.
-(2018), and SARFA(Puri et al.
-2020) for the PPO agent(Schulman et al.
-2017; Nguyen 2020) trained in the Super Mario Bros.(Kauten 2018) environment.
-Saliency level ranges from blue(low) to red(high).
-Row 1: The chosen action is to move left.
-The explanation is that Mario is attempting to avoid the Piranha Plants on top of the green pipes.
-VisionMask accurately identified the two Piranha Plants on the pipes.
-While RISE and Greydanus highlighted almost the entire image and LIME and SARFA highlighted areas not directly related to the explanation.
-Row 2:The chosen action is to jump right with acceleration.
-The explanation is that Mario intends to land on top of the Koopas to clear his path and get reward.
-Again VisionMask clearly highlighted the two Koopas whereas the other saliency maps failed to identify the correct region.
-Deep Reinforcement Learning (DRL) is a powerful technology in machine intelligence, widely used for many applications, including ChatGPT (Brown et al.
-2020; Achiam et al.
-2023) and Alpha Go Zero (Silver et al.
-2017).
-It enables effective end-to-end training of decision models that map an agent’s observations to actions, aiming to maximizing the long term expected rewards.
-However, understanding a DRL agent’s decision-making process is challenging, due to the inherent lack of explainability in the high-dimensional, non-linear structure of its underlying Deep Neural Network (DNN).
-The lack of transparency undermines human users’ trust and acceptance, particularly in safety-critical scenarios, such as medical diagnosis and military operations.
-To ensure wide acceptance and efficient use of this technology, it is crucial to provide accurate and human-interpretable explanations.
-These considerations motivated the recent interests in Explainable Reinforcement Learning (XRL).
-Some existing works integrate explanation generation into the agent model and train them together.
-We refer to these as model-based approaches.
-For example, Interpretable DQN (i-DQN) (Annasamy and Sycara 2019) implements a deep Q-network with an attention mechanism, using the attention weights between state and action embeddings to represent their importance.
-Saliency-guided Q-networks (SGQNs) (Bertoin et al.
-2022) calculate the importance of inputs as a derivative of the Q-function and train the critic model to both minimize the value prediction error and maximize the fidelity of the explanation.
-Some other works moved away from DNN and adopt special policy architectures that are inherently more interpretable (Liu et al.
-2018; Verma et al.
-2018; Bastani, Pu, and Solar-Lezama 2018).
-However, these models usually experience performance loss compared to the DNN-based policy model due to the limitations imposed by the special structure.
-Those model-based approaches cannot provide explanations for a trained DRL agent, unless they retrain the agent using their specific models, which is prohibitively expensive.
-And the retrain process may compromise the agent’s performance and integrity as it strives to maintain a balance between the optimality and explainability.
-One of the most widely used techniques to provide explanations for machine learning models is the attribution-based method, which assigns credits to input features based on their contribution to AI agent’s decisions.
-This technique was originally developed to explain image classification models in computer vision.
-It generates saliency maps by perturbing the input image to reveal their contributions to the classification process without modifying the classifier itself (Dabkowski and Gal 2017; Petsiuk, Das, and Saenko 2018; Ribeiro, Singh, and Guestrin 2016).
-Inspired by this, several works have explored to use attribution-based technique to explain DRL models while treating the agent as a black-box.
+Deep Reinforcement Learning (DRL) is a powerful technology in machine intelligence, widely used for many applications Sutton et al.
+(1999).
+However, understanding a DRL agent’s decision-making process is challenging, due to the inherent lack of explainability in the high-dimensional, non-linear structure of its underlying Deep Neural Network (DNN) Heuillet et al.
+(2021); Hickling et al.
+(2023).
+The lack of transparency undermines users’ trust, driving the development of Explainable AI (xAI).
+Various methods in computer vision have been proposed to enhance the transparency of AI systems Ribeiro et al.
+(2016); Selvaraju et al.
+(2017); Bach et al.
+(2015); Shrikumar et al.
+(2017); Lundberg and Lee (2017).
+At the core, they share a common foundation: attributing the classifier’s outputs to more interpretable features and using a saliency map to visualize these attributions.
+Their only differences are how these attributions are calculated.
+A high-quality attribution-based explanation should meet several key criteria.
+First, it should demonstrate faithfulness, meaning that including features with high attribution should lead the model to the target output, and excluding them should prevent it.
+Second, it should exhibit specificity, ensuring that only critical features receive high attribution.
+Sometime this is also referred to as sparseness.
+Finally, it should be robust, meaning the explanation should remain consistent and not change significantly with minor variations in the input.
+Attribution-based explanation has been studied for DRL models.
 Greydanus et al.
-(2018) perturb the input features using Gaussian blur at fixed intervals and measure how the perturbation changes the action probability.
-The difference in action probability before and after perturbation are used to generate a saliency-map for the importance of each input feature.
-Similarly, Iyer et al.
 (2018) and  Puri et al.
-(2020) compute the importance score based on the differences in Q values before and after the perturbation.
-The latter also improve the accuracy of the saliency map by normalizing the effects of perturbation between target and non-target actions.
-The perturbation-based methods do not alter the agent model, nor do they impose any specific requirements on it.
-Hence they are applicable to any trained DRL agents.
-However, these techniques have some significant limitations.
-First, to manage complexity, perturbations are performed according to a set of predefined rules, such as using Gaussian blur or occluding with a constant value, which severely limits the search space they can explore.
-For DRL agents that map visual input directly to actions, due to the large input dimensions, there is no guarantee that the predefined perturbation can find the set of pixels that trigger the agent decision.
-Secondly, perturbation-based methods do not learn from past experiences, nor do they improve performance by comparing similarities and differences across different input states and actions.
-Thirdly, each input feature (e.g., input pixel) is perturbed independently, with no guarantee that the joint state of the perturbed pixels represents meaningful objects.
-As shown in Table 1, the saliency maps generated by the perturbation-based method either consider all pixels equally important or identify regions in the image that do not contain meaning objects.
-To address the aforementioned challenges, we propose VisionMask, a model trained end-to-end to generate attribution-based explanations for vision-based DRL without perturbation or altering the agent.
-VisionMask ouputs a saliency map that highlights the critical regions in the agent’s visual input responsible for triggering the agent’s action.
-The saliency map reveals the importance of each pixel in the decision making process, and will also be referred to as importance map or importance mask in the rest of the paper.
-VisionMask is trained using self-supervised contrastive learning.
-No human generated label is required.
-The training of VisionMask does not alter the agent model, thereby preserving the agent’s performance and integrity.
-The (fixed) policy model of the DRL agent is integrated into the training pipeline to generate the loss, and involved in the backpropagation to pass the gradient to the mask generation model.
-Therefore, each VisionMask model is trained to explain a specific DRL agent.
-The training process is agent agnostic, meaning that it works with any agent with a policy model that maps visual input to actions.
-VisionMasks’s learnable capability allows it to accumulate knowledge from previous experiences to improve future explanations.
-During testing time, VisionMask does not need to access agent’s internal state to generate the importance mask.
-The mask is generated by a forward pass of the VisionMask model.
-As shown in Table 1, VisionMask accurately identifies the critical regions in the visual input that explains the agent’s action.
-Beyond providing precise explanations directly from the visual input, VisionMask also facilitates counterfactual analysis.
-By selectively activating or deactivating different critical regions, meaningful counterfactual scenarios can be generated and assessed.
-This is similar to the perturbation-based approach, however, at a much higher level.
+(2020) utilized policy distributional shifts as the basis for attribution in RL.
+Specifically, they calculate attribution of a feature as the difference in Q/V values or action distributions between the original and perturbed states.
+For example, given agent policy $\pi$, the attribution of a feature is proportional to $E_{s^{\prime}}(|\pi(s)-\pi(s^{\prime})|_{2})$ where $s$ stands for the original state and $s^{\prime}$ represents perturbed states generated for this feature.
+By calculating the attributions for all features, a saliency map $m$ can be created.
+However, the perturbation-based explanations lack faithfulness.
+Since each perturbation focuses only on local features while ignoring the joint impact of feature combinations, overlaying the saliency map with the original state, $(m\odot s)$, does not result in a feature combination that leads the agent to the target action distribution $\pi(s)$.
+A better approach to enhance faithfulness is to learn a model to predict the saliency map $m$ that minimizes the difference between $\pi(m\odot s)$ and $\pi(s)$.
+Explainer Stalder et al.
+(2022) leveraged this idea by training an explanation model for an image classifier.
+However, Explainer categorizes class labels into target and non-target for each training sample and focus on learning saliency map (or mask) only for the target label while treat all non-target labels as a single group.
+Unlike a (well trained) image classifier, where predictions for non-target labels are typically close to 0, DRL agent in many scenarios, does not exhibit a clear preference for the actions.
+Non-target actions may sometimes have probabilities only slightly lower than those of target actions.
+Analyzing how masking the feature may affect the non-target action probability provides additional information that can be used to train the explanation model more effectively.
+The above analysis motivates us to design a trainable saliency map generator for attribution-based explanations and train it using two channels of contrastive information:(i) Action-wise contrast: We believe that environment states contain features that motivate the DRL agent to select both target action and non-target actions.
+However, the target action is ultimately chosen because it corresponds to higher reward or has a stronger presence.
+For each action, a saliency map can be generated as an explanation.
+Choosing features according to the saliency map for a non-target action should push the agent away from the target action, and vice versa.
+This inspired us to treat the saliency map of the non-target action $m_{a^{\prime}}$ and the target action $m_{a}$ as a negative pair, which can be leveraged for contrastive learning Chopra et al.
+(2005); Schroff et al.
+(2015); Gutmann and Hyvärinen (2010).
+(ii) Feature-wise contrast: To exclude irrelevant features (e.g.
+background) from the saliency map, explanations also need to be discriminative in filtering out such information.
+When only irrelevant features are accessible to the agent, the resulting action distribution should be as uniform as possible.
+Therefore, the target action’s saliency map ($m$) and its inverted counterpart ($\tilde{m}=1-m$) form another negative pair for contrastive learning.
+In this work, we present VisionMask as an RL explainer that is contrastively trained to generate saliency maps to explain agent’s actions.
+We specifically focus on agents that maps images to actions and consider each pixel value as the interpretable input feature, although the similar technique could be extended to other type of features.
+We carefully design the objective function to enable self-supervised contrastive learning of explanations from both action-wise and feature-wise perspectives, fostering the generation of more faithful explanations.
+We conduct evaluation on six RL environments with five baselines based on faithfulness, robustness, and sparseness.
+Quantitatively, VisionMask outperforms the baselines in terms of faithfulness, while exhibiting strong robustness and high sparseness.
+Qualitatively, we compared VisionMask with the baselines in two settings: visual comparison and human studies.
+In the visual comparison, VisionMask provides sharper explanations that align more closely with human interpretations, as demonstrated by counterfactual examples.
+In the human studies, VisionMask’s explanations help users better understand the agent’s decisions and calibrate appropriate trust.
 The contribution of our work can be summarized as the following:
-  • We present VisionMask, a novel attribution method for explainable visual reinforcement learning that generates action-specific importance masks.
-Generated using a single forward pass, the mask does not only highlight the regions in the visual input that explain the agent’s behavior but also facilitates counterfactual analysis.
-  • VisionMask is model-agnostic in that it works with any vision-based DRL agent without modifying the agent model.
-  • VisionMask is trained using self-supervised contrastive learning. No human intervention is needed for data labeling.
-  • The fidelity of the explanation is measured by the agreement between the agent’s actions generated from the original visual input versus from the masked visual input.
-We quantify this agreement using accuracy, precision, recall and F1-score.
-In average, VisionMask outperforms other attribution-based explanation models by 30.08% on F1-score and 14.9% on insertion accuracy..
+  • We present VisionMask, a novel attribution method for explainable visual reinforcement learning that generates action-specific saliency maps.
+To the best of our knowledge, this is the first work that learns saliency map in a self-supervised contrastive learning manner, yielding highly faithful explanations across various RL environments.
+  • VisionMask is model-agnostic in that it works with any vision-based DRL agent and requires only the agent’s input state and output action during inference time.
+VisionMask does not make any modification on the agent, however, each VisionMask model is trained specifically for a given agent by including the fixed agent in the back-propagation path for gradient descent.
+  • We conducted extensive experiments, both qualitatively and quantitatively, to demonstrate the improvements of the saliency map in terms of faithfulness, robustness, and sparseness, as well as a comprehensive ablation study for each component.
 2 Background and Related Work
 The nested structure and non-linear operation of DNN make it challenging for humans to understand how the outputs are derived from inputs.
-Some existing works address this challenge by explaining model outputs through input attributions (Dabkowski and Gal 2017; Petsiuk, Das, and Saenko 2018; Ribeiro, Singh, and Guestrin 2016; Fong and Vedaldi 2017).
-For example, Randomized Input Sampling for Explanation of Black-box Models (RISE) (Petsiuk, Das, and Saenko 2018) is a perturbation-based approach that explains neural image classifiers by applying randomly generated masks to the input image and assessing their impact.
-Local Interpretable Model-agnostic Explanations (LIME)(Ribeiro, Singh, and Guestrin 2016) tries to explain local instances by approximating them within a nearby vicinity using a linear decision model, where the explainable elements are superpixels (i.e., small image regions of similar pixels).
-Built on top of the DNNs, the lack of transparency of the DRL agents (Samek, Wiegand, and Müller 2017; Ras et al.
-2022; Angelov and Soares 2020) undermines humans trust and hinders their adoption.
+Some existing works address this challenge by explaining model outputs through input attributions Petsiuk et al.
+(2018); Ribeiro et al.
+(2016); Fong and Vedaldi (2017).
+For example, Randomized Input Sampling for Explanation of Black-box Models (RISE) Petsiuk et al.
+(2018) is a perturbation-based approach that explains neural image classifiers by applying randomly generated masks to the input image and assessing their impact.
+Local Interpretable Model-agnostic Explanations (LIME)Ribeiro et al.
+(2016) tries to explain local instances by approximating them within a nearby vicinity using a linear decision model, where the explainable elements are superpixels (i.e., small image regions of similar pixels).
+Built on top of the DNNs, the lack of transparency of the DRL agents undermines humans trust and hinders their adoption.
 Several studies have addressed this challenging problems.
-According to (Qing et al.
-2022), existing efforts can be categorized into four main approaches: model explaining, reward explaining, task explaining, and state explaining.
-Model explaining relies on inherently interpretable model architectures or auxiliary reasoning mechanisms to generate explanations (Topin et al.
-2021; Jiang and Luo 2019; Landajuela et al.
-2021).
+According to Qing et al.
+(2022), existing efforts can be categorized into four main approaches: model explaining, reward explaining, task explaining, and state explaining.
+Model explaining relies on inherently interpretable model architectures or auxiliary reasoning mechanisms to generate explanations Topin et al.
+(2021).
 However, these self-explanatory or symbolic models often suffer from decreased performance compared to state-of-the-art neural network based RL policies and may lack the representational power needed to learn more complex policies.
-Reward explanation typically involves using explicitly designed explainable reward functions to generate explanations (Juozapaitis et al.
-2019; Ashwood, Jha, and Pillow 2022).
-Task explanation considers a policy as a composition of sub-task and action sequences, explaining the behavior in terms of relationships among sub-tasks (Shu, Xiong, and Socher 2017; Sodhani, Zhang, and Pineau 2021).
+Reward explanation typically involves using explicitly designed explainable reward functions to generate explanations Ashwood et al.
+(2022).
+Task explanation considers a policy as a composition of sub-task and action sequences, explaining the behavior in terms of relationships among sub-tasks Shu et al.
+(2017).
 These approaches often assume the existence of a reward decomposition scheme or a predefined sub-task partition mechanism, which may not hold true in all RL environments.
@@
 State explaining methods can further be divided into three categories: attention-based, Shapley value-based and perturbation-based mechanisms, which are detailed below.
-Attention is a common approach used for explainable RL in several existing works (Peng, Riedl, and Ammanabrolu 2022; Xu et al.
-2020; Annasamy and Sycara 2019).
-However, similar to i-DQN (Annasamy and Sycara 2019), these methods cannot explain a given DRL model because their attention model must be trained alongside the agent model.
-Shapley value (Shapley et al.
-1953), a concept from game theory, has also been introduced into XRL.
-SVERL (Beechey, Smith, and Şimşek 2023) provides a theoretical analysis on how Shapley values can be applied to explain value and policy network in DRL.
+Attention is a common approach used for explainable RL in several existing works Annasamy and Sycara (2019).
+However, similar to i-DQN Annasamy and Sycara (2019), these methods cannot explain a given DRL model because their attention model must be trained alongside the agent model.
+Shapley value Shapley and others (1953), a concept from game theory, has also been introduced into XRL.
+SVERL Beechey et al.
+(2023) provides a theoretical analysis on how Shapley values can be applied to explain value and policy network in DRL.
 However, this approach is still in its early stages and assumes prior knowledge of the transition model of the environment, which is not realistic for realistic applications.
-Perturbation-based methods (Greydanus et al.
-2018; Iyer et al.
-2018; Puri et al.
-2020) compute saliency maps of the input features by comparing the action probability or value function before and after perturbation.
+Perturbation-based methods Greydanus et al.
+(2018); Iyer et al.
+(2018); Puri et al.
+(2020) compute saliency maps of the input features by comparing the action probability or value function before and after perturbation.
 Greydanus et al.
@@
 Furthermore, there is no guarantee that the perturbed input remains physically meaningful.
+Figure 1: Architecture of VisionMask. State-action pairs $(s,a)$ are sampled from $\mathcal{D}_{E}$.
 3 VisionMask
-This section presents VisionMask, an innovative agent-agnostic framework for XRL.
-The primary goal is to produce action-wise importance-mask that attribute explainable elements in the current state to each possible action.
-The definition of explainable elements may vary across different applications.
-However, in this paper, we focus on identifying explanatory pixel regions in the agent’s visual input.
-Figure 1: Architecture of VisionMask. In this example $a=1$ so $m_{1}$ is the active mask $m_{a}$.
+In this section, we present our VisionMask architecture.
+The primary goal is to generate action-wise saliency maps that attribute the most relevant features in the state to each action.
+For agents that map images to actions, the features and states correspond to pixels and images.
 3.1 Problem Formulation
-We approach the task of VisionMask as an imitation learning problem based on Behavioral Cloning (BC) (Bain and Sammut 1995; Ross, Gordon, and Bagnell 2011).
-Formally, we define the environment as a Markov Decision Process (MDP) $\{\mathcal{S},\mathcal{A},P,R,\gamma\}$, where $\mathcal{S}$ represents the state space; $\mathcal{A}$ denotes the action space; the state transition function $P:\mathcal{S}\times\mathcal{A}\to\Delta(\mathcal{S})$ depicts the transition between states based on actions, where $\Delta(\mathcal{S})$ represents the set of probability distributions over states; the reward function $r:\mathcal{S}\times\mathcal{A}\to\mathbb{R}$ provides the immediate reward for state-action pairs; and $\gamma\in[0,1]$ and $\pi:\mathcal{S}\to\Delta(\mathcal{A})$ represent the discount factor and policy.
+Formally, we define the environment as a Markov Decision Process (MDP) $\{\mathcal{S},\mathcal{A},P,R,\gamma\}$, where $\mathcal{S}$ represents the state space; $\mathcal{A}$ denotes the action space with $|\mathcal{A}|=K$; the state transition function $P:\mathcal{S}\times\mathcal{A}\to\Delta(\mathcal{S})$ depicts the transition between states based on actions, where $\Delta(\mathcal{S})$ represents the set of probability distributions over states; the reward function $R:\mathcal{S}\times\mathcal{A}\to\mathbb{R}$ provides the immediate reward for state-action pairs; $\gamma\in[0,1]$ is discount factor and $\pi:\mathcal{S}\to\Delta(\mathcal{A})$ represent policy.
 Return $G$ is defined as $G=\sum_{k=0}^{\infty}\gamma^{k}R_{k+1}$, and the expected cumulative reward of a policy $\pi$ is $\mathbb{E}_{\pi}[G]=\mathbb{E}_{\pi}[\sum_{k=0}^{\infty}\gamma^{k}R_{k+1}]$, where the expectation is taken with respect to the initial state distribution, transition probabilities, and action probabilities determined by $\pi$.
-VisionMask operates on a given pre-trained expert policy $\pi_{E}$ such that
-$\pi_{E}\approx\pi^{*}=\operatorname*{arg\,max}_{\pi}\mathbb{E}_{\pi}[G],$
-where $\pi^{*}$ is the optimal policy.
-We can obtain a dataset of expert demonstrations $\mathcal{D}_{E}=\{(s_{i},\pi_{E}(s_{i}))\}_{i=1}^{N}$ consisting of $N$ state-action pairs, from trajectories sampled while executing $\pi_{E}$ in the environment.
-Our goal is to learn an explainer $f_{\theta^{*}}$ that minimizes the loss
-$\theta^{*}=\operatorname*{arg\,min}_{\theta}\sum_{(s,a)\in\mathcal{D}_{E}}L(a,s,\theta),$
-where $L$ is the training loss function to be discussed in section 3.3.
-The explainer function $f_{\theta}:\mathcal{S}\to[0,1]^{K\times d_{s}}$ (with $d_{s}$ representing the state dimension and $K$ the number of candidate actions) provides the attributions of each action to each feature in the state $s$.
+VisionMask operates on a given trained expert policy $\pi_{E}$ such that $\pi_{E}\approx\pi^{*}=\arg\max_{\pi}\mathbb{E}_{\pi}[G],$ where $\pi^{*}$ is the optimal policy.
+We can obtain a dataset of expert demonstrations $\mathcal{D}_{E}=\{(s_{i},a_{i}\sim\pi_{E}(s_{i}))\}_{i=1}^{N}$ consisting of $N$ state-action pairs, from trajectories sampled while executing $\pi_{E}$ in the environment.
+Our goal is to learn an explainer $f_{\theta^{*}}$ that minimizes the loss $\theta^{*}=\arg\min_{\theta}\sum_{(s,a)\in\mathcal{D}_{E}}\mathcal{L}(a,s,\theta)$ where $\mathcal{L}$ is the training loss function to be discussed in section 3.3.
+The explainer function $f_{\theta}:\mathcal{S}\to[0,1]^{K\times d_{s}}$, where $d_{s}$ represents the feature size of the state $s\in\mathcal{S}$ and $K$ denotes the number of candidate actions, predicts the attributions of each action to each feature in the state $s$.
+The value of the output is bounded within the range $[0,1]$, with the $(i,j)$th element indicating the $j$th feature’s attribution to the $i$th action.
 In the case of visual DRL, $s$ is the visual input of the agent and $d_{s}$ is the number of pixels in $s$, i.e., $d_{s}=W\times{H}$, with $W$ and $H$ representing the width and height of the visual input.
-The value of each element in the attribution is bounded within the range $[0,1]$, with the $(i,j)$th element indicates the $j$th pixel’s attribution to the $i$th action.
-This attribution will then be overlaid onto the visual input, masking unimportant pixels while retaining important ones.
-The overlaid visual input will be analyzed by the policy model again.
-We will update parameters $\theta$ of explainer $f$ so that the overlaid visual input always produces the same action as the original one.
-The expert policy $\pi_{E}$ remains frozen during the training process of $f_{\theta}$, meaning its parameters are fixed.
+The output of the explainer can be viewed as $K$ saliency maps, $M=\{m_{0},m_{1},....,m_{K-1}\}$, each corresponds to an action.
+A saliency value $m_{i}[x,y]$ near 1 indicates that the pixel $(x,y)$ has significant contributions to action $i$, whereas a value near 0 denotes irrelevance.
+Overlaying the $i$th saliency map with the state $s$ will highlight the input features that lead the agent to the $i$th action.
 3.2 Architecture
-As shown in Fig.
-1, the expert dataset $\mathcal{D}_{E}$ is collected using the expert policy $\pi_{E}$.
-From this dataset $\mathcal{D}_{E}$, state-action pairs $(s_{i},a_{i})$ are sampled.
-Subsequently, the explainer $f_{\theta}$ takes the state $s_{i}$ as input and produces a set of action-wise masks $M=\{m_{0},m_{1},....,m_{K-1}\}$, where $K$ denotes the number of candidate actions and $m_{i}\in{[0,1]^{W\times{H}}}$.
-A mask value $m_{i}[x,y]$ near 1 indicates that the pixel $(x,y)$ has significant contributions to the corresponding action $i$, whereas a value near 0 denotes irrelavance.
-This process shares similarities with image segmentation, where each pixel is attributed to an object or background.
-Hence, we structure the explainer $f_{\theta}$ akin to the widely used image segmentation model, DeepLabV3 (Chen et al.
-2017), but retrain it with the expert demonstrations $\pi_{E}$.
-To make sure that the output masks are bounded to the range $[0,1]$, a sigmoid function is applied at the output of $f_{\theta}$
-Following (Stalder et al.
-2022), we partition the masks in $M$ into two sets: the active set and the inactive set.
-Unlike (Stalder et al.
-2022), our active set has only one mask, $m_{a}$ where $a$ is the action selected by expert policy $\pi_{E}$ll the other masks, $m_{i}$, where $i\neq{a}$, belong to the inactive set.
-Each mask $m_{i}$ contains the attributions for the current agent state that explain the corresponding action $i$.
-The explainer will be trained to produce masks for pixels that contribute the most to the policy’s action selection.
-We aggregate the masks in the inactive set into a non-target mask $n$ by taking the maximum value pixel-wise.
-We also calculate a complement mask $\tilde{m_{a}}=\mathbf{1}-m_{a}$ representing the less important regions for the selected action $a$.
-Then the masks $m_{a}$ and $\tilde{m_{a}}$ are overlaid onto the original state $s$ to generate two masked states $s_{m}$ and $s_{\tilde{m}}$ using the following overlay function:
-$s_{m}=s\odot m_{a}+r\odot(\tilde{m_{a}}),s_{\tilde{m}}=s\odot\tilde{m_{a}}+r\odot(m_{a})$
+As shown in Figure 1, we first collect the expert dataset $\mathcal{D}_{E}$ using the expert policy $\pi_{E}$.
+From this dataset $\mathcal{D}_{E}$, state-action pairs $(s_{i},a_{i})$ are sampled and fed to VisionMas $f_{\theta}$ to generate the set of saliency maps $M$.
+Generating the saliency map from given visual input is a dense prediction task that shares similarities with image segmentation, where each pixel is assigned a value to indicate whether it belongs to an object or background.
+Hence, we structure the explainer $f_{\theta}$ akin to the widely used image segmentation model, DeepLabv3 Chen et al.
+(2017), however, retrain it using self-supervised contrastive learning.
+To make sure that the output saliency value are bounded to the range $[0,1]$, a sigmoid function is applied at the output of $f_{\theta}$.
+For each $m_{i}\in M$, we also calculate a complement map $\tilde{m}_{i}=\mathbf{1}-m_{i}$ highlighting the irrelevant regions for the action $i$.
+Then the masks $m_{i}$ and $\tilde{m}_{i}$ are overlaid onto the original state $s$ to generate two masked states $s_{i}$ and $\tilde{s}_{i}$ using the following overlay function:
+$s_{i}=s\odot m_{i}+r\odot(\tilde{m}_{i}),\tilde{s}_{i}=s\odot\tilde{m}_{i}+r\odot(m_{i})$ (1)
 where $\odot$ is Hadamard Product and $r$ is a reference value.
-The final step is to feed these masked states $s_{m}$ and $s_{\tilde{m}}$ through the expert policy $\pi_{E}$ to generates action distributions $\mathbf{p}$ and $\mathbf{\tilde{p}}$, respectively:
-$\mathbf{p}=\pi_{E}(\mathbf{s}_{\mathbf{m}}),\mathbf{\tilde{p}}=\pi_{E}(\mathbf{s}_{\mathbf{\tilde{m}}}).$
-The model $f_{\theta}$ is optimized so that $\mathbf{p}$ matches the action distribution chosen for the unmasked state $s$, while $\mathbf{\tilde{p}}$ does not.
-It is worth mentioning that the agent may be sensitive to the reference value in the overlay function.
 Numerous options exist for the reference value, such as setting the pixel to zero, assigning a constant value, blurring the pixel, or cropping it.
-In this study, we opt for a constant value as the reference value, based on the sensitivity of the agent to zero and blurring, as indicated by experimental results.
-We use different reference values for different environments (See Appendix 4).
+Empirical study shows that setting the reference to the background gives the best results.
+More details can be found in the ablation study 4.4.
+See Appendix F for the background values for different environments.
+To generate self-supervised contrastive loss to train the model, we query the agent to obtain the corresponding logits $z_{i}=z(s_{i})$ and $\tilde{z}_{i}=z(\tilde{s}_{i})$, where $z_{i},\tilde{z}_{i}\in\mathbb{R}^{K}$ , and the action probability distributions $p_{i}=\text{softmax}(z_{i})$ and $\tilde{p}_{i}=\text{softmax}(\tilde{z}_{i})$, where $p,p_{i}\in[0,1]^{K}$, $0\leq{i}<K$.
+By concatenating each $p_{i}$ and $\tilde{p}_{i}$, we have the the action probability distributions of each mask $\mathbf{p},\mathbf{\tilde{p}}\in\mathbb{R}^{K\times K}$,
+$\mathbf{p}=[p_{1},p_{2},\dots,p_{k}]^{T}\quad\mathbf{\tilde{p}}=[\tilde{p}_{1},\tilde{p}_{2},\dots,\tilde{p}_{k}]^{T}$
 3.3 Training Loss
-To train the explainer to capture the most important regions in the state, we carefully designed the training loss function $\mathcal{L}$ as follows:
-$\mathcal{L}(\mathbf{s},\mathbf{a},\theta,\mathbf{n})$ $=\mathcal{L}_{bc}(\mathbf{s},\mathbf{a})+\lambda_{e}\mathcal{L}_{e}(\mathbf{s},\mathbf{a})+\lambda_{avg}\mathcal{L}_{avg}(\mathbf{s},\mathbf{a},\mathbf{n})$
-$+\lambda_{\text{smooth}}\mathcal{L}_{\text{smooth}}(\mathbf{s},\mathbf{a},\mathbf{n})+\lambda_{L2}\lVert\theta\rVert^{2}_{2},$
-where $\mathcal{L}_{bc}(\mathbf{s},\mathbf{a})$ is Behavior Cloning loss, $\mathcal{L}_{e}(\mathbf{s},\mathbf{a})$ is entropy loss, $\mathcal{L}_{avg}(\mathbf{s},\mathbf{a},\mathbf{n})$ is the mask average loss, $\mathcal{L}_{\text{smooth}}(\mathbf{s},\mathbf{a},\mathbf{n})$ is the smoothness loss, $\lVert\theta\rVert^{2}_{2}$ is $L_{2}$ regularization and $\lambda_{e},\lambda_{avg},\lambda_{\text{smooth}},\lambda_{L2}$ are regularization hyper-parameters.
-Behavioral Cloning loss $\mathcal{L}_{bc}$.
-Our goal is to train the explainer to select regions in the visual input that that are sufficient and necessary for the agent to choose action $\mathbf{a}$.
-Hence, we design the behavior cloning loss as cross entropy loss between the action distribution $\mathbf{p}$ predicted for the overlaid state $s_{m}$ and the target $\mathbf{a}$ chosen for the original state $s$:
-$\mathcal{L}_{bc}(\mathbf{s},\mathbf{a})=-\frac{1}{K}\sum_{k=1}^{K}\llbracket k=\mathbf{a}\rrbracket\log(\mathbf{p}[k]).$
+To enable the agent contrastively learn the saliency map $m_{a}$, we carefully designed the training loss function $\mathcal{L}$ as follows:
+$\mathcal{L}(\mathbf{s},\mathbf{a},\theta)=\mathcal{L}_{a}(\mathbf{s},\mathbf{a})+\lambda_{ne}\mathcal{L}_{ne}(\mathbf{s})+\lambda_{area}\mathcal{L}_{area}(\mathbf{s},\mathbf{a})$
+where $\mathcal{L}_{a}(\mathbf{s},\mathbf{a})$ is action-wise contrastive action loss, $\mathcal{L}_{ne}(\mathbf{s},\mathbf{a})$ is feature-wise loss, $\mathcal{L}_{area}(\mathbf{s},\mathbf{a},\mathbf{n})$ is the area size loss and $\lambda_{ne},\lambda_{area}$ are regularization hyper-parameters.
+Action-wise contrastive loss $\mathcal{L}_{a}$.
+Let $a$ denote the target action chosen by the agent.
+Our primary goal is to learn explanations faithful to the $a$, making $(a,p_{a})$ the only positive pair.
+Furthermore, the explanation must be discriminative, meaning it should clearly distinguish the target action $a$ from all other possible actions.
+As a result, every other pair of $(a,p_{i\neq a})$ is treated as a negative pair.
+Then, we compute the cross-entropy loss between these pairs,
+$\mathcal{L}_{pos}(\mathbf{s},\mathbf{a})=-\frac{1}{K}\sum_{k=1}^{K}\llbracket k=\mathbf{a}\rrbracket\log(\frac{\text{exp}(z_{aa})}{\sum_{i=1}^{K}\text{exp}(z_{ai})})$ (2)
+$\mathcal{L}_{neg}(\mathbf{s},\mathbf{a})=-\frac{1}{K}\sum_{k=1}^{K}\llbracket k=\mathbf{a}\rrbracket\log(\frac{\text{exp}(z_{aa})}{\sum^{K}_{i=1}\text{exp}(z_{ia})})$ (3)
+where $z_{ij}$ is the logit for action $j$ when querying the policy with the masked state $s_{i}$.
+The contrastive action loss $\mathcal{L}_{a}=\mathcal{L}_{pos}+\mathcal{L}_{neg}$.
 Here $\llbracket k=\mathbf{a}\rrbracket$ denotes the indicator function which returns 1 if $k$ is the same as label $\mathbf{a}$, and 0 otherwise.
-Negative Entropy loss $\mathcal{L}_{e}$.
-To ensure that the visual input regions selected by $m_{a}$ is sufficient and necessary for the agent to make decisions, we also need to make sure that the unselected region, i.e., $s_{\tilde{m_{a}}}$, does not provide useful information for action selection, hence the action distribution $\mathbf{\tilde{p}}$ should follow a uniform distribution.
+Note that we do not compute the loss with $\mathbf{p}_{k,i\neq a}$, as ensuring the faithfulness of the target action $a$ is our primary objective here.
+Feature-wise loss $\mathcal{L}_{ne}$.
+To ensure that the visual input regions selected by $m_{i}$ is necessary for the agent to make decisions, we also need to make sure that the unselected region, i.e., $s_{\tilde{m_{i}}}$, does not provide useful information for action selection, hence the action distribution $\mathbf{\tilde{p}}$ should follow a uniform distribution.
 Motivated by this rationale, we define negative entropy loss regarding $\tilde{\mathbf{m}}$ as the following:
-$\mathcal{L}_{e}(\mathbf{s},\mathbf{a})=\frac{1}{K}\sum^{K}_{k=1}\tilde{\mathbf{p}}[k]\log\tilde{\mathbf{p}}[k].$
-Mask average loss $\mathcal{L}_{avg}$.
-A low effort way to minimize $\mathcal{L}_{bc}$ and $\mathcal{L}_{e}$ is to include all pixels in the importance mask, $m_{a}$, and no pixel in the complement mask, $\tilde{m_{a}}$, which obviously is not a valid solution.
+$\mathcal{L}_{ne}(\mathbf{s})=\frac{1}{K^{2}}\sum_{ij}\tilde{\mathbf{p}}_{ij}\log\tilde{\mathbf{p}}_{ij}.$ (4)
+Area size loss $\mathcal{L}_{area}$.
+A low effort way to minimize $\mathcal{L}_{a}$ and $\mathcal{L}_{ne}$ is to include all pixels in the mask, $m_{i}$, and no pixel in the complement mask, $\tilde{m_{i}}$, which obviously is not a valid solution.
 We need to ensure that each importance mask only consists a small number of crucial pixels.
-Thus, we define $\mathcal{L}_{avg}$ as follows:
-$\mathcal{L}_{avg}(\mathbf{s},\mathbf{a},\mathbf{n})=\frac{1}{W\times H}\sum_{i,j}\mathbf{m}[i,j]+\frac{1}{W\times H}\sum_{i,j}\mathbf{n}[i,j],$
-where, $\mathbf{n}$ represents the non-target mask aggregated from the inactive set. By minimizing the mask average loss, we limit the number of pixels selected by the masks.
-Smoothness loss $\mathcal{L}_{\text{smooth}}$. Finally, we minimize the Variation between adjacent pixels to generate smoother masks:
-$\mathcal{TV}(x)$ $=\frac{1}{W\times H}\sum_{i,j}(|x[i,j]-x[i+1,j]|$
-$+|x[i,j]-x[i,j+1]|)$
-$\mathcal{L}_{\text{smooth}}(\mathbf{s},\mathbf{a},\mathbf{n})=\mathcal{TV}(m)+\mathcal{TV}(n)$
-The overall architecture of VisionMask is given in Fig. 1. The loss will backpropagate through the frozen expert model $\pi_{E}$ to train the explainer model $f_{\theta}$.
+Thus, we define $\mathcal{L}_{area}$ using L1 norm as follows:
+$\mathcal{L}_{area}(\mathbf{s})=\frac{1}{K}\sum_{k}(|\frac{1}{Z}\sum_{i,j}m_{k}[i,j]-a_{max}|)$ (5)
+where $Z$ it the number of pixels in state.
+SMB Enduro Seaquest
+Method Acc.
+Del.
+$\downarrow$ Ins.
+$\uparrow$ LLE $\downarrow$ Sp.
+$\uparrow$ Acc.
+Del.
+$\downarrow$ Ins.
+$\uparrow$ LLE $\downarrow$ Sp.
+$\uparrow$ Acc.
+Del.
+$\downarrow$ Ins.
+$\uparrow$ LLE $\downarrow$ Sp.
+$\uparrow$
+LIME 78.7 27.0 30.2 69.4 98.3 88.1 36.3 38.0 47.9 90.2 90.1 14.3 21.4 17.3 97.7
+RISE 80.5 18.9 38.0 1.9 2.0 90.3 37.6 39.1 0.01 0.5 92.0 8.1 29.0 0.02 1.4
+Greydanus 16.9 56.4 20.2 20.8 63.3 27.4 34.6 33.3 0.25 82.9 44.6 14.4 7.88 0.12 75.1
+SARFA 16.9 55.6 21.4 68.7 65.5 27.9 33.3 33.4 0.26 77.0 44.8 7.6 8.7 0.13 75.2
+Explainer 92.2 23.6 49.0 38.1 81.2 90.2 34.2 33.1 0.22 81.2 95.3 13.4 30.0 0.9 97.5
+VisionMask 95.9 20.4 67.6 38.0 82.3 98.7 32.9 41.2 0.5 80.0 99.6 6.4 34.3 0.9 97.6
+MsPacman VizDoom Highway
+Method Acc.
+Del.
+$\downarrow$ Ins.
+$\uparrow$ LLE $\downarrow$ Sp.
+$\uparrow$ Acc.
+Del.
+$\downarrow$ Ins.
+$\uparrow$ LLE $\downarrow$ Sp.
+$\uparrow$ Acc.
+Del.
+$\downarrow$ Ins.
+$\uparrow$ LLE $\downarrow$ Sp.
+$\uparrow$
+LIME 84.5 32.7 37.7 9.1 96.6 72.1 35.1 11.2 22.1 96.4 77.8 20.0 23.64 52.7 80.5
+RISE 92.6 26.2 45.7 0.01 0.59 83.8 14.6 14.5 0.02 0.5 82.5 20.0 22.1 0.05 0.6
+Greydanus 46.1 39.1 14.2 0.17 57.4 75.2 20.7 17.2 0.12 62.1 92.7 20.9 24.7 0.93 83.8
+SARFA 42.4 18.8 15.6 0.13 66.5 76.1 16.7 17.4 0.55 66.6 97.6 20.4 25.6 0.43 83.0
+Explainer 97.6 19.6 59.2 0.22 81.2 84.2 14.4 17.6 0.2 65.5 95.0 20.37 24.4 1.5 83.9
+VisionMask 98.7 17.0 62.8 0.2 89.5 87.8 14.2 18.1 3.89 47.8 98.1 19.8 25.8 1.14 84.8
+Table 1: Quantitative results on SMB, Enduro, Seaquest, MsPacman, VizDoom and Highway of VisionMask against 5 baselines.
+Five metrics are compared.
+The faithfulness is measured by Accuracy(Acc.), Deletion(Del.) and Insertion(Ins.) metrics(%); Robustness is measured by Local Lipschitz Estimate (LLE)(%); And Complexity is measured by Sparseness(Sp.)(%).
+Blue represents second best results.
+Figure 2: Qualitative examples of VisionMask and five baselines across three environments.
+(a)–(c) show the saliency map overlaid on the input image and counterfactuals where regions are removed based on VisionMask’s map.
+(a) Human: ”Mario moves left to avoid the Piranha Plant.” VisionMask correctly highlights the Plant; removing it changes the action from ’Left’ to ’Jump right’.
+(b) Human: ”The agent moves left at constant speed due to the front car.” VisionMask identifies both; removing the car allows acceleration instead of ’Right + accelerate’.
+(c) Human: ”The agent fires because a shark follows.” VisionMask detects the shark and oxygen bar; removing the shark stops firing.
+Additional examples in Appendix H.
 4 Experiments
-Env Metric VisionMask LIME RISE Greydanus SARFA
-Mario Acc. $\mathbf{0.9076}\pm\scriptstyle{0.0124}$ $0.7868\pm\scriptstyle{0.0027}$ $0.8053\pm\scriptstyle{0.0118}$ 0.1692 0.1692
-Precision $\mathbf{0.7109}\pm\scriptstyle{0.0189}$ $0.2448\pm\scriptstyle{0.0200}$ $0.3364\pm\scriptstyle{0.0046}$ 0.6154 0.6154
-Recall $\mathbf{0.7005}\pm\scriptstyle{0.0020}$ $0.2467\pm\scriptstyle{0.0042}$ $0.3245\pm\scriptstyle{0.0069}$ 0.1752 0.1800
-F1-score $\mathbf{0.6995}\pm\scriptstyle{0.0058}$ $0.2277\pm\scriptstyle{0.0056}$ $0.3046\pm\scriptstyle{0.0082}$ 0.2764 0.2753
-Enduro Acc. $\mathbf{0.9061}\pm\scriptstyle{0.0077}$ $0.8811\pm\scriptstyle{0.0087}$ $0.9033\pm\scriptstyle{0.0047}$ 0.2744 0.2790
-Precision $0.6058\pm\scriptstyle{0.0084}$ $0.4167\pm\scriptstyle{0.0062}$ $0.5363\pm\scriptstyle{0.0054}$ $\mathbf{0.6449}$ 0.6383
-Recall $\mathbf{0.5535}\pm\scriptstyle{0.0049}$ $0.4551\pm\scriptstyle{0.0046}$ $0.5245\pm\scriptstyle{0.0072}$ 0.0839 0.0889
-F1-score $\mathbf{0.5060}\pm\scriptstyle{0.0082}$ $0.3809\pm\scriptstyle{0.0018}$ $0.4738\pm\scriptstyle{0.0032}$ 0.1283 0.1418
-Seaquest Acc. $\mathbf{0.9531}\pm\scriptstyle{0.0030}$ $0.9094\pm\scriptstyle{0.0079}$ $0.9201\pm\scriptstyle{0.0051}$ 0.4463 0.4476
-Precision $0.5809\pm\scriptstyle{0.0022}$ $0.2647\pm\scriptstyle{0.0066}$ $0.4631\pm\scriptstyle{0.004}$ 0.8083 $\mathbf{0.8371}$
-Recall $\mathbf{0.4880}\pm\scriptstyle{0.0013}$ $0.2423\pm\scriptstyle{0.0082}$ $0.2830\pm\scriptstyle{0.0047}$ 0.0945 0.0983
-F1-score $\mathbf{0.4714}\pm\scriptstyle{0.0050}$ $0.2073\pm\scriptstyle{0.01443}$ $0.2733\pm\scriptstyle{0.0110}$ 0.1640 0.1701
-MsPacman Acc. $\mathbf{0.9768}\pm\scriptstyle{0.0177}$ $0.8451\pm\scriptstyle{0.0294}$ $0.9264\pm\scriptstyle{0.0220}$ 0.4616 0.4235
-Precision $0.7671\pm\scriptstyle{0.0160}$ $0.1878\pm\scriptstyle{0.0298}$ $0.5590\pm\scriptstyle{0.0097}$ 0.8255 $\mathbf{0.9102}$
-Recall $\mathbf{0.8658}\pm\scriptstyle{0.0039}$ $0.1726\pm\scriptstyle{0.0036}$ $0.5587\pm\scriptstyle{0.0035}$ 0.2189 0.1913
-F1-score $\mathbf{0.7699}\pm\scriptstyle{0.0096}$ $0.1597\pm\scriptstyle{0.0086}$ $0.5122\pm\scriptstyle{0.0078}$ 0.3307 0.3067
-Table 2: Fidelity scores on four environments and confusion matrices in the Mario environment. We run each experiment 3 times with different random seeds and report the mean and standard deviation.
-We evaluated VisionMask across four different environments and compared its performance against four baseline methods.
+In this section, we begin by outlining the experimental setup.
+We then present quantitative and qualitative analyses to evaluate our approach.
+Additionally, we provide counterfactual explanations to demo VisionMask’s faithfulness and sensitivity.
+Finally, we perform an ablation study to assess the contribution of each component.
+4.1 Experimental Setup
 Environment Selection.
-Four agent-centered RL environments were selected to showcase the effectiveness of VisionMask: Super Mario Bros (SMB)(Kauten 2018), and the Atari games Seaquest, Enduro, and MsPacman.
-We select these environments following existing works (Puri et al.
-2020; Greydanus et al.
-2018).
+We conduct experiments across three types of environments: Super Mario Bros (SMB) Kauten (2018), Enduro, Seaquest and MsPacman Bellemare et al.
+(2013); Machado et al.
+(2018) for 2D game, Highway-env Leurent (2018) for autonomous driving simulation and VizDoom Wydmuch et al.
+(2018) for 3D game.
 Baseline Selection.
-We compared VisionMask’s performance against four perturbation-based explainer, LIME (Ribeiro, Singh, and Guestrin 2016) RISE (Petsiuk, Das, and Saenko 2018), Greydanus et al.
-(2018) and SARFA(Puri et al.
-2020).
-For LIME, we used the original public implementation with the default configuration available from (Ribeiro 2017).
-For RISE, we adopt the public implementation from torchray (Fong, Patrick, and Vedaldi 2019).
-For Greydanus et al.
-(2018) and SARFA, we used the respective public implementation provided by the authors.
-To ensure a fair comparison, we did not include any explainers that require retraining the agent model.
+We mainly compare our model with perturbation-based baselines for black-box RL such as Greydanus Greydanus et al.
+(2018) and SARFA Puri et al.
+(2020).
+In addition, we also compared with three techniques originally designed to explain image classifiers, including a learning-based method, Explainer Stalder et al.
+(2022), and two perturbation-based methods, LIME Ribeiro et al.
+(2016) and RISE Petsiuk et al.
+(2018).
+Although these methods focus on image classification, their main concept is similar to ours: to attribute a given label (action) to a subset of visual features.
+We use the public implementation from torchray Fong et al.
+(2019) for RISE and the original published implementations for other baselines.
+Among all the baselines, the Explainer is the most similar to VisionMask, as both are learning-based approaches.
+However, VisionMask employs action-contrastive learning and is trained using different regularization.
+In the experimental results, we demonstrate that these differences significantly enhances VisionMask’s performance.
 Expert Policy.
-For reproducibility, we did not train the expert policies ourselves; instead, we used publicly available implementations.
-In the SMB environment, We apply the trained PPO (Schulman et al.
-2017) policy from (Nguyen 2020).
-In the Atari environments, we used three trained PPO policies from Stable-Baselines3 (Raffin et al.
-2021) .
-4.1 Assessment of Explanation Fidelity
-We quantify fidelity using four metrics: accuracy, precision, recall, and F1-score.
-Fidelity measures the agreement between the agent’s action selected for the original visual input and the action selected for the masked visual input.
-In this context, we define the accuracy as the following:
-$\text{Accuracy}=\frac{1}{|\mathcal{D}_{\text{test}}|}\sum_{i=1}^{|\mathcal{D}_{\text{test}}|}\llbracket\pi_{E}(s_{i})=\pi_{E}(s_{i}^{\prime})\rrbracket,$
+We use the open-source pre-trained PPO Schulman et al.
+(2017) agents from Nguyen (2020) for SMB environment and from stable-baselines3 Raffin et al.
+(2021) for Enduro, Seaquest and MsPacman environments.
+We train the DQN Mnih et al.
+(2013) agents from scratch for the VizDoom and Highway-env environments.
+See Appendix A for agents details.
+Dataset.
+We collect state-action pairs for each environment and split the data into 80% for training, 10% for validation, and 10% for testing.
+All results in this section are reported using the test split.
+We make this dataset publicly available; see Appendix B for details.
+4.2 Quantitative Analysis
+Metrics.
+Since there is no ground truth explanations Adebayo et al.
+(2020), it is crucial to select appropriate metrics to evaluate the trustworthiness of explanations.
+Designing robust and consistent metrics for XAI remains an unresolved challenge Hedström et al.
+(2023).
+To alleviate the inconsistency between metrics and avoid selection bias, we closely follow the LATEC Klein et al.
+(2024) benchmark to evaluate performance across three dimensions: Faithfulness, Robustness, and Complexity.
 […]
-[…]
-Values of these fidelity metrics are shown in Table 2.
-VisionMask consistently scores higher than LIME and RISE.
-Greydanus and SARFA have a higher precision than VisionMask, however, a much lower recall and F1-score.
-This is probably because they included almost all pixels from the visual input in the importance map.
-They also have a lower accuracy compared to VisionMask.
-Refer to caption Refer to caption
-Original state VisionMask
-Refer to caption Refer to caption
-Insertion curve Deletion curve
-Figure 2: Examples of insertion and deletion AUC curve
-Env Method 0.25 0.5 1.0
-$\downarrow$ del $\uparrow$ ins $\downarrow$ del $\uparrow$ ins $\downarrow$ del $\uparrow$ ins
-Mario VisionMask 0.2044 ±0.1891 0.4156 ±0.2937 0.1813 ±0.1730 0.4142 ±0.2788 0.1714 ±0.1577 0.4216 ±0.2622
-LIME 0.3447 ±0.2664 0.2314 ±0.2189 0.3231 ±0.2568 0.2446 ±0.2232 0.2693 ±0.1958 0.3024 ±0.2107
-RISE 0.2750 ±0.2510 0.2940 ±0.2624 0.2331 ±0.2326 0.3317 ±0.2761 0.1892 ±0.1897 0.3796 ±0.2678
-Greydanus 0.5708 ±0.2109 0.1963 ±0.1658 0.5671 ±0.2146 0.1952 ±0.1650 0.5634 ±0.2169 0.2018 ±0.1691
-SARFA 0.5579 ±0.2223 0.2071 ±0.1756 0.5565 ±0.2235 0.2117 ±0.18 0.5556 ±0.2246 0.2143 ±0.1866
-Enduro VisionMask 0.3693 ±0.2854 0.3952 ±0.3004 0.3589 ±0.2911 0.3979 ±0.2995 0.3491 ±0.2948 0.4021 ±0.2941
-LIME 0.3780 ±0.2886 0.3636 ±0.3025 0.3723 ±0.2914 0.3680 ±0.3002 0.3629 ±0.2937 0.3799 ±0.2883
-RISE 0.4081 ±0.2818 0.3585 ±0.3066 0.3970 ±0.2861 0.3714 ±0.2938 0.3759 ±0.2912 0.3906 ±0.2785
-Greydanus 0.3762 ±0.2988 0.3312 ±0.3194 0.3626 ±0.3018 0.3317 ±0.3192 0.3457 ±0.3075 0.3329 ±0.3188
-SARFA 0.3472 ±0.3060 0.3327 ±0.3188 0.3378 ±0.3114 0.3336 ±0.3186 0.3331 ±0.3157 0.3340 ±0.3186
-Seaquest VisionMask 0.0721 ±0.0291 0.3256 ±0.2596 0.0684 ±0.0289 0.3331 ±0.2650 0.0652 ±0.0268 0.3430 ±0.2635
-LIME 0.1982 ±0.1871 0.1397 ±0.1662 0.1909 ±0.1865 0.1425 ±0.1672 0.1434 ±0.1214 0.2141 ±0.1812
-RISE 0.1198 ±0.0797 0.2422 ±0.2560 0.0987 ±0.0658 0.2638 ±0.2654 0.0809 ±0.0505 0.2896 ±0.2660
-Greydanus 0.2939 ±0.2381 0.0658 ±0.0345 0.2165 ±0.1834 0.0708 ±0.0542 0.1439 ±0.0980 0.0788 ±0.0825
-SARFA 0.1133 ±0.0732 0.0819 ±0.0978 0.0888 ±0.0418 0.0851 ±0.1104 0.0764 ±0.0307 0.0869 ±0.1177
-MsPacman VisionMask 0.2251 ±0.1051 0.6350 ±0.3056 0.2037 ±0.1255 0.6231 ±0.3337 0.1956 ±0.1217 0.6278 ±0.2778
-LIME 0.4831 ±0.2700 0.1901 ±0.1708 0.4348 ±0.2455 0.2346 ±0.1875 0.3265 ±0.1801 0.3771 ±0.1890
-RISE 0.4270 ±0.3140 0.2299 ±0.1220 0.3512 ±0.2995 0.2937 ±0.1325 0.2621 ±0.2273 0.4577 ±0.1394
-Greydanus 0.5669 ±0.2476 0.1285 ±0.0646 0.4903 ±0.2518 0.1336 ±0.0843 0.3905 ±0.2281 0.1417 ±0.1029
-SARFA 0.3511 ±0.2528 0.1463 ±0.1168 0.2524 ±0.1573 0.1531 ±0.1328 0.1880 ±0.0873 0.1561 ±0.1425
-Table 3: Insertion and deletion metrics for fractions 0.25, 0.5, and 1.0.
-The symbol $\uparrow$ (resp.
-$\downarrow$) indicates that higher (resp.
-lower) values are better.
-We report the mean $\pm$ standard deviation over testing samples.
-4.2 Assessment of Effectiveness
-An effective explainer should precisely identify the exact input regions that contribute to the expert decision, without including irrelevant areas.
-To quantify the effectiveness of explanations, We adopt the insertion and deletion metrics from RISE (Petsiuk, Das, and Saenko 2018).
-The rationale behind these metrics is as follows: If certain pixels in the visual state are crucial to the action (i.e., the true cause of the action), deleting them from the state or inserting them into the background should result in a corresponding decrease or increase in the probability of the specific action.
-By plotting the relationship between the probability of the target action and the number of pixels added or removed, we obtain an insertion curve and a deletion curve.
-The Area Under the Curve (AUC) of those plots represents the effectiveness of the explanation.
-A higher AUC for the insertion curve and a lower AUC for the deletion curve stand for more effective explanation.
-The order of insertion and deletion affects the result.
-We sort the pixels based on their values in the importance map and insert or delete them in descending order of their importance.
-Figure 2 shows an example of insertion and deletion curves with corresponding AUC values displayed at the top right corner.
-See Appendix D for the pseudocode of each metric.
-It is worth noting that the effectiveness varies as different number of pixels are inserted or deleted from the original input.
-Instead of showing the actual curve for each explainer under different games, we report the AUC under 25%, 50%, and 100% level of insertion and deletion.
-The results are given in Table 3.
-VisionMask outperforms all other methods across all environments, with an average insertion score that is 14.9% higher and a deletion score that is 8.8% lower compared to the other methods.
-4.3 Counterfactual Explanations
-Original state VisionMask Counterfactual
-[Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
-Move left Jump right
-[Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
-Left Left + acceleration
-[Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
-Down Fire Down right
-Figure 3: Counterfactual analysis with VisionMask.
-By removing the most important regions identified by VisionMask from the visual input, we can create scenarios for counterfacutal analysis.
-Table 3 provides three examples of this approach.
-In the first example, VisionMask correctly identified the Piranha Plants as the reason the agent chose to move left.
-After removing the Piranha Plant by replacing the highlighted pixels with reference value and run the modified image by the agent again, we observed that the agent’s action changed to jumping right, as the threat from the Piranha Plant no longer existed.
-In the second example, VisionMask correctly identified the edge of the road and a car in front as the reason why the agent chose to move left (with current speed).
-After we removed the car in front the agent chose to move left with acceleration.
-In the third example, VisionMask identified the shark in the bottom left as the cause of the downward and fire action.
-When the corresponding shark was removed, the agentno longer chose to fire.
-Appendix I, J and K contain more examples.
-We believe such explanations can further help users to understand the RL agent’s behavior.
+Insertion and deletion measure the impact on the target action probability by progressively adding or removing pixels from the original input based on the descending order of their attribution scores, with the most important pixels being inserted/deleted first.
+The inserted/deleted input image will be processed by the expert agent again to obtain the probability of the target action.
+By plotting the probability against the number of pixels added or removed, we obtain an insertion and a deletion curve.
+The insertion/deletion performance is measured by the area under the curve (AUC) of the insertion/deletion curves.
+A larger (smaller) AUC for the insertion (deletion) curve means that including (removing) the important pixels identified by the explainer can effectively increase (reduce) the probability of the target action.
+Hence a larger (smaller) AUC of insertion (deletion) curve indicates more accurate attribution prediction.
+The pseudo-code of the detailed information of deletion and insertion could be found in Appedix E.
+For robustness, we report the Local Lipschitz Estimate (LLE) scores Alvarez Melis and Jaakkola (2018), which quantify the local smoothness of explanations by estimating the Lipschitz constant within a specific neighborhood.
+The Lipschitz constant measures the maximum rate of change of the function, ensuring explanation does not vary too rapidly within the state’s neighborhood.
+Given state $s_{i}$ and neighborhood size $\epsilon$, LLE defined as
+$\hat{L}(s_{i})=\underset{s_{j}\in\mathcal{N}_{\epsilon}(s_{i})}{\text{argmax}}\frac{\|f_{\theta}(s_{i})-f_{\theta}(s_{j})\|_{2}}{\|s_{i}-s_{j}\|_{2}}$
+where $\mathcal{N}_{\epsilon}(s)=\{s^{\prime}\in X\mid\|s-s^{\prime}\|\leq\epsilon\}$.
+We evaluate complexity with Sparseness Chalasani et al.
+(2020), which uses the Gini index on the vector of absolute attribution values sorted in non-descending order.
+Sparseness ensures that features genuinely influencing the output have substantial contributions, while insignificant or only slightly relevant features should have minimal contributions.
+A higher Sparseness indicates more contrastive attribution values and hence more understandable explanations  Chalasani et al.
+(2020).
+Results.
+In Table 1, our VisionMask achieves the best performance in terms of faithfulness (i.e., Acc, Del, Ins) in all testing environment except SMB, where its deletion score is slightly lower than RISE.
+However, RISE has significantly lower insertion score and accuracy in this environment.
+Hence, our explanations are more aligned with the agent’s decision-making process.
+Moreover, compared to, Explainer Stalder et al.
+(2022), VisionMask exhibits much higher faithfulness, which suggests the effectiveness of action contrastive learning.
+Overall, learning-based model performs better compared to perturbation based approach.
+In terms of robustness, RISE achieves the best performance in terms of LLE score across all settings.
+This is because, as a perturbation-based method, RISE uses the same type of perturbation that we used to generate the neighbor image for LLE score calculation.
+Hence the perturbation almost has no impact to it.
+On the other hand, RISE has the lowest sparsity, which means the attribution predicted by RISE are evenly distributed and contains little information.
+For Complexity, benefiting from the binary mask of LIME which is much more sparse compared to other baselines, LIME has the highest sparseness score.
+However, it has the worst performance in terms of LLE score.
+Overall, VisionMask achieves the best balance between the robustness and sparseness.
+See Appendix for the Radar map.
+Metrics Stalder L1
+Acc. 95.6 95.9
+Ins. $\uparrow$ 64.3 67.6
+Del. $\downarrow$ 21.1 20.4
+LLE $\downarrow$ 56.6 62.6
+Sp. $\uparrow$ 73.8 62.3
+Hours 0.72 0.68
+Speedup - $1.1\times$
+Sp./LLE 1.3 1.0
+(a) Area size regularization. L1 is better for faithfulness and faster. Time is estimated by training 10 epochs.
+Metrics FCN UNet DeepLabv3
+Acc. 94.2 94.7 96.0
+Ins. $\uparrow$ 56.5 56.8 67.5
+Del. $\downarrow$ 19.6 19.7 20.4
+LLE $\downarrow$ 100.0 86.3 62.6
+Sp. $\uparrow$ 58.0 19.8 62.3
+Sp./LLE 0.58 0.23 0.99
+(b) $f_{\theta}$ model. DeepLabv3 achieves better results.
+Metrics $0.1$ $0.2$ $0.4$
+Acc. 95.9 94.5 95.8
+Ins. $\uparrow$ 57.0 56.9 57.0
+Del. $\downarrow$ 19.5 19.5 19.5
+LLE $\downarrow$ 96.0 93.8 99.5
+Sp. $\uparrow$ 79.0 58.6 48.1
+Sp./LLE 0.82 0.62 0.48
+(c) Max area size $a_{\text{max}}$. Maximum area size of $0.1$ produces favorable results.
+Metrics background black mean blur
+Acc. 97.8 95.9 97.6 94.4
+Ins. $\uparrow$ 67.5 51.7 54.5 56.9
+Del. $\downarrow$ 20.4 16.8 18.2 25.0
+LLE $\downarrow$ 62.6 54.4 92.6 99.6
+Sp. $\uparrow$ 62.3 28.5 45.5 56.6
+Sp./LLE 1.0 0.52 0.49 0.57
+(d) Reference value. Using background produce better faithfulness values.
+Comp. C. +NE +NE+L1 +NE+L1+TV
+Acc. 96.7 96.5 95.9 95.6
+Ins. $\uparrow$ 56.7 56.5 67.6 64.6
+Del. $\downarrow$ 23.7 23.8 20.4 20.7
+LLE $\downarrow$ 99.3 96.0 62.6 68.8
+Sp. $\uparrow$ 24.0 23.8 62.3 60.5
+Sp./LLE 0.24 0.25 0.99 0.88
+(e) Component. Negative Entropy (NE) with L1 area regularization performs better in terms of faithufulness. Contrastive-only (C.), Stalder Area Loss (S.), and Total Variation (TV.)
+Table 2: VisionMask ablation experiments on SMB. Default settings are marked in gray.
+4.3 Qualitative Analysis
+In this section, we conduct two qualitative analyses across two settings, visual comparision and human studies.
+Visual Comparision.
+We present example explanations from three environments, SMB, Enduro, and Seaquest, in Figure 2, along with some counterfactual analysis generated from the explanations provided by VisionMask.
+The examples show that both LIME and RISE fail to generate interpretable explanations.
+LIME’s superpixels are too large to capture the specific regions, while RISE’s explanations include almost all pixels.
+Explainer generates more accurate and interpretable explanations compared to LIME and RISE.
+This suggests that purely perturbation-based approaches may fail in RL due to the high dynamics of environments and the lack of learning ability.
+Greydanus et al.
+(2018) and SARFA generate better explanations than LIME and RISE but often focus on irrelavent objects or background.
+In contrast, VisionMask accurately highlights the relevant regions, providing sharp explanations that are both accurate and interpretable.
+By removing some of the regions highlighted by the VisionMask, counterfactual analysis could be performed to answer questions like: ”Why this action instead of that one?”
+Figure 3: Human studies. Comparison of total responses and positive responses across questions.
+Human Studies.
+To assess whether the saliency maps improve humans’ understanding of the agent’s decisions, we conduct human studies for SMB, Enduro and Seaquest.
+We present 63 participants with 9 state-action pairs accompanied by the saliency maps generated by VisionMask.
+Participants are then asked whether the saliency maps help them better understand the decision-making process of the RL agents.
+We record the total number of responses as well as the number of positive responses, as shown in Figure 3.
+Over 89% of participants find the saliency maps generated by VisionMask helpful in understanding the agents’ decisions.
+Details about this study can be found in the Appendix D.
+4.4 Ablation Studies
+Table 2 shows the results for ablation studies. In all tables, the column marked in gray represents the default setting of VisionMask used to generate results in previous sections.
+Area size regularization.
+In this experiment, we replace the default L1 area regularization with the Min-Max Area Loss from Explainer Stalder et al.
+(2022).
+Without this constraint, Explainer tends to produce masks where all values are one in order to achieve the best performance.
+Explainer first vectorizes and sorts the mask elements, and then applies penalties to pixels falling outside the range $[a,b]$, where the hyperparameters $a$ and $b$ specify the minimum and maximum allowable area sizes.
+Instead, VisionMask uses L1 area regularization to constrain the maximal area, as we do not wish to impose constraints on the minimum area size.
+To implement action contrastive learning, we need to generate masks for actions that are not selected, and these masks should be allowed to be all zeros if necessary.
+As shown in Table 2a, using L1 area regularization improves the faithfulness of the explanation and increases training speed.
+Max area size.
+We further vary the maximum permitted area size $a_{\text{max}}$ to evaluate its impact.
+As shown in Table 2c, reducing $a_{\text{max}}$ significantly increases the sparseness as it constrains the model to focus on the most discriminative regions without compromising other performance metrics.
+The $0.1$ setting achieves the best results, as expected.
+$\mathbf{f_{\theta}}$ model.
+In Table 2b, we compare the performance of three different versions of VisionMask with three different segmentation models: our default DeepLabv3 Chen et al.
+(2017), Fully Convolutional Network (FCN)Long et al.
+(2015), and UNetRonneberger et al.
+(2015).
+The results show that DeepLabv3 achieves a better performance.
+Reference Value.
+We evaluate the impact of hyperparameter $r$ in Equation 1 by setting the reference value to background, black, mean, and blur.
+The ”black” reference sets $r$ to 0, the ”mean” reference sets $r$ to the RGB mean value of the entire dataset, and the ”blur” reference applies Gaussian blur (kernel is 39, $\sigma=15$) to the image and use the blurred pixel value as $r$.
+Table 2d compares the performance of VisionMask with these four different reference values.
+We choose the background as reference value as it provides a more balanced performance across all metrics.
+Details about the background reference value are provided in Appendix F.
+Loss Function Components.
+In Table 2e, we evaluate the impact of different components in the loss function on the performance of VisionMask.
+The results show that, although adding negative entropy and L1 area regularization slightly reduces accuracy, it improves the insertion and deletion scores and significantly boosts the Sp./LLE ratio.
+In contrast, adding total variation to produce smoother masks has minimal impact on overall performance.
 5 Conclusion
@@
 VisionMask generates explanations with higher fidelity and better effectiveness compared to existing attribute-based methods.
-It is our future plan to extend this approach to multi-modality input and couple the visual explanation generated by the VisionMake with other information such as agent’s long-term goals and future rewards.
-Acknowledgements
-This research is partially supported by the Air Force Office of Scientific Research (AFOSR), under contract FA9550-24-1-0078.
-The paper was received and approved for public release by AFRL on May 28th 2024, case number AFRL-2024-2908 .
+It is our future plan to extend this approach to multi-modality input and couple the visual explanation generated by the VisionMask with other information such as agent’s long-term goals and future rewards.
+Acknowledgments
+This research is partially supported by the Air Force Office of Scientific Research (AFOSR), under contract FA9550-24-1-0078, and NSF award CNS-2148253.
+The paper was received and approved for public release by Air Force Research Laboratory (AFRL) on May 28th 2024, case number AFRL-2024-2908 .
 Any opinions, findings, and conclusions or recommendations expressed in this material are those of the authors and do not necessarily reflect the views of AFRL or its contractors.
-Appendix A Technical Appendix
-Appendix B Meanings of actions across four environments
-In Table 4, we present the list of actions and their corresponding meanings for the four environments evaluated in the experiment section.
-Value Meaning Value Meaning
+Appendix A Agent details
+config value
+learning rate 0.001
+batch size 32
+$\epsilon$ 0.5
+$\gamma$ 0.99
+buffer size 50000
+target update interval 1000
+Table 3: VizDoom DQN
+config value
+learning rate 0.0005
+batch size 32
+$\epsilon$ 0.5
+$\gamma$ 0.8
+buffer size 15000
+target update interval 50
+Table 4: Highway-env DQN
+As mentioned in the section 4, we adopt the pretrained PPO models from stable-baselines3 Raffin et al.
+[2021] for Enduro, Seaquest and MsPacman.
+Check the Hugging Face repositories for detailed information on the PPO agents for Enduro [^1] https://huggingface.co/sb3/ppo-EnduroNoFrameskip-v4, Seaquest [^2] https://huggingface.co/sb3/ppo-SeaquestNoFrameskip-v4, and MsPacman [^3] https://huggingface.co/sb3/ppo-MsPacmanNoFrameskip-v4.
+For SMB, we use SMB-PPONguyen [2020] [^4] https://github.com/vietnh1009/Super-mario-bros-PPO-pytorch.
+Inputs for above agents are 4 continuous RGB frames without frame skipping of shape $84\times 84$.
+For VizDoom, we train the DQN agent from scratch.
+The agent consists of four convolutional layers, each with a kernel size of 3, stride of 1, and filter sizes ranging from 64 to 512 (64, 128, 256, 512).
+Inputs are 5 greyscale frames wihout frame skipping of shape $60\times 40$.
+The training setting is in Table 3.
+For Highway-env, we train the DQN agent of 2 convolutional layers (kernel size 3, stride 1 and filter size 32).
+Inputs are 4 greyscale frames without frame skipping of shape $128\times 64$.
+Details of training is in Table 4.
+Appendix B Datasets
+env train val test state dim state format actions
+SMB 7754 970 969 $4\times 256\times 240$ RGB 7
+Enduro 8000 1000 1000 $4\times 160\times 210$ RGB 9
+Seaquest 8000 1000 1000 $4\times 160\times 210$ RGB 9
+MsPacman 8000 1000 1000 $4\times 160\times 210$ RGB 9
+VizDoom 13336 1667 1667 $5\times 60\times 40$ Grey 7
+Highway-env 15940 1992 1993 $4\times 64\times 128$ Grey 5
+Table 5: Datasets details of six environments
+With the aforementioned agents, we collect total state-action pairs of 9693 for SMB in 31 different worlds and stages following Nguyen [2020], 10000 for Enduro, Seaquest and MsPacman, 16670 for VizDoom and 19925 for Highway-env.
+See Table 5 for details.
+The datasets is publicly available  [^5] https://huggingface.co/datasets/0pow0/visionmask_datasets
+Mario Meaning Enduro Meaning
 0 Stay in place 0 No operation
@@
 8 Accelerate and move left
-(a) Mario environment (b) Atari Enduro environment
-Value Meaning Value Meaning
+Seaquest Meaning MsPacman Meaning
 0 No operation 0 No operation
@@
 17 Fire and move down-left
-(c) Atari Seaquest environment (d) Atari MsPacman environment
-Table 4: Action meanings for SMB (Nguyen 2020), Atari Enduro, Seaquest and MsPacman (Brockman et al. 2016)
-Appendix C Reference values used across different environments
-In our experiment, we found that the PPO agents we adopted from Stable Baselines3 (Raffin et al.
-2021) and Mario-PPO (Nguyen 2020) are sensitive to reference values.
-We tried using a constant value (zero) and blur as reference values but did not obtain clear and reasonable importance masks.
-However, using reference values that are visually similar to the environment background gave us better results.
-In Figure 4, we list the reference values we used in each environment.
-Specifically, for the SMB and Atari Enduro environments, we use the background color of the content in each state as the reference value.
-Note that the reference value for different states may vary, as the background content of each state can be different.
-For the Atari Seaquest environment, since all the states share the same background, we use a single reference value for each state, which combines the empty sea background and the oxygen bar background.
-For the Ms.
-Pacman environment, we simply used a constant value of the background color blue as the reference value.
-Figure 4: Reference values for four environment evaluated.
+VizDoom Meaning Highway-env Meaning
+0 TURN LEFT 0 LANE LEFT
+1 TURN RIGHT 1 IDLE
+2 ATTACK 2 LANE RIGHT
+3 MOVE LEFT 3 FASTER
+4 MOVE RIGHT 4 SLOWER
+5 MOVE FORWARD
+6 MOVE BACKWARD
+Table 6: Action meanings for SMB, Enduro, Seaquest and MsPacman
+Appendix C Actions
+In Table 6, we present the list of actions and their corresponding meanings for the six environments evaluated in the experiment section.
+Appendix D Example of the human studies
+Figure 4: Example interface for our human studies
+In Figure 4, we present a screenshot of the interface used for our human studies.
+We selected three examples from SMB, Enduro, and Seaquest.
+Participants were asked to provide Yes/No responses based on whether the saliency map’s explanation aligned with their understanding and if it was reasonable.
+Appendix E Algorithms for insertion and deletion metrics
+Algorithms 1-2 present the pseudocode for evaluating the insertion and deletion metrics of Petsiuk et al. [2018].
+Algorithm 1 Evaluate insertion metric
+1:Input: target mask $m$, state $s$, action $a$, expert policy $\pi_{E}$, fraction of pixels inserted $\alpha$, Number of pixels NN
+2:Initialize $m^{\prime}\leftarrow\text{zeros}(m)$, $p\leftarrow\emptyset$
+[…]
+4:for $k\leftarrow 0$ to $\lfloor\alpha N\rfloor-1$ do
+5:  $(x,y)\leftarrow\mathcal{I}[k]$, $m^{\prime}[x,y]\leftarrow 1.0$
+6:  $p\leftarrow p\cup\pi_{E}(s\odot m^{\prime}+r\odot(1-m^{\prime}))$
+7:end for
+8:return $\text{AUC}(p)$
+Algorithm 2 Evaluate deletion metric
+1:Input: target mask $m$, state $s$, action $a$, expert policy $\pi_{E}$, fraction of pixels deleted $\alpha$, Number of pixels NN
+2:Initialize $m^{\prime}\leftarrow\text{ones}(m)$, $p\leftarrow\emptyset$
+[…]
+4:for $k\leftarrow 0$ to $\lfloor\alpha N\rfloor-1$ do
+5:  $(x,y)\leftarrow\mathcal{I}[k]$, $m^{\prime}[x,y]\leftarrow 0.0$
+6:  $p\leftarrow p\cup\pi_{E}(s\odot m^{\prime}+r\odot(1-m^{\prime}))$
+7:end for
+8:return $\text{AUC}(p)$
+Figure 5: Reference values for four environment.
 The content column represents a sub-figure of the original state that contains the objects and background.
@@
 MsPacman [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
-Appendix D Algorithms for insertion and deletion metrics
-Algorithms 1-2 present the pseudocode for evaluating the insertion and deletion metrics of (Petsiuk, Das, and Saenko 2018).
-Algorithm 1 Evaluate insertion metric
-1:Input: target mask $m$, state $s$, action $a$, expert policy $\pi_{E}$, fraction of pixels inserted $\alpha$, Number of pixels $N$
-2:Initialize $m^{\prime}\leftarrow\text{zeros}(m)$, $p\leftarrow\emptyset$
-3:$\mathcal{I}\leftarrow\text{indices}(\text{sort}(\text{flatten}(m)))$
-4:for $k\leftarrow 0$ to $\lfloor\alpha N\rfloor-1$ do
-5:     $(x,y)\leftarrow\mathcal{I}[k]$, $m^{\prime}[x,y]\leftarrow 1.0$
-6:     $p\leftarrow p\cup\pi_{E}(s\odot m^{\prime}+r\odot(1-m^{\prime}))$
-7:end for
-8:return $\text{AUC}(p)$
-Algorithm 2 Evaluate deletion metric
-1:Input: target mask $m$, state $s$, action $a$, expert policy $\pi_{E}$, fraction of pixels deleted $\alpha$, Number of pixels $N$
-2:Initialize $m^{\prime}\leftarrow\text{ones}(m)$, $p\leftarrow\emptyset$
-3:$\mathcal{I}\leftarrow\text{indices}(\text{sort}(\text{flatten}(m)))$
-4:for $k\leftarrow 0$ to $\lfloor\alpha N\rfloor-1$ do
-5:     $(x,y)\leftarrow\mathcal{I}[k]$, $m^{\prime}[x,y]\leftarrow 0.0$
-6:     $p\leftarrow p\cup\pi_{E}(s\odot m^{\prime}+r\odot(1-m^{\prime}))$
-7:end for
-8:return $\text{AUC}(p)$
-Appendix E Object importance percentage calculation
-[…]
-Appendix F Time efficiency
-To assess the time efficiency of mask generation, we compare the total time taken for mask generation using VisionMask against that of two baseline methods on testing datasets from four environments.
-Note that the time we measure does not include the training time for VisionMask.
-We can see that LIME and RISE are slower than VisionMask.
-All the experiments are conducted on A40 GPU.
-Env VisionMask LIME RISE
-SMB 00:00:10 25:45:02 00:51:41
-Enduro 00:00:10 1:05:19 00:08:14
-Seaquest 00:00:12 1:34:49 00:10:58
-MsPacman 00:00:10 3:02:28 00:08:59
-Table 5: Time taken to generate importance masks for three methods across four environments.
-The time format is hh:mm:ss.
-The testing dataset for SMB contains 970 states, while Atari Enduro, Seaquest, and Ms.
-Pacman contain 4000 states each.
-Appendix G Experimental details
-Dataset details
-Recall from Section 3.1 that we collected the expert dataset $\pi_{E}$ with the agent being explained as expert in four envrionments.
-In the SMB environment(Kauten 2018), $\pi_{E}^{\text{Mario}}$ comprises a total of 9,692 state-action pairs.
-In Atari Enduro, Seaquest and MsPacman (Brockman et al.
-2016) environment, $\pi_{E}^{\text{Atari}}$ includes 10,000 total state-action pairs.
-We utilize a 80%-10%-10% train-valid-test split on all expert dataset $\pi_{E}$.
-Pre-processing data
-We transform the RGB input to greyscale with the following function then resize it to $84\times 84$:
-1def rgb_to_grayscale(img, num_output_channels: int = 1):
-2 if img.ndim < 3:
-3 raise TypeError(f"Input image tensor should have at least 3 dimensions, but found {img.ndim}")
-4 # _assert_channels(img, [1, 3])
-5
-6 if num_output_channels not in (1, 3):
-7 raise ValueError("num_output_channels should be either 1 or 3")
-8
-9 if img.shape[-3] == 3:
-10 r, g, b = img.unbind(dim=-3)
-11 # This implementation closely follows the TF one:
-12 # https://github.com/tensorflow/tensorflow/blob/v2.3.0/tensorflow/python/ops/image_ops_impl.py#L2105-L2138
-13 l_img = (0.299 * r + 0.587 * g + 0.114 * b).to(img.dtype)
-14 l_img = l_img.unsqueeze(dim=-3)
-15 else:
-16 l_img = img.clone()
-17
-18 if num_output_channels == 3:
-19 return l_img.expand(img.shape)
-20
-21 return l_img
-Model details
-We utilize the segmentation model DeeplabV3 as the $f_{\theta}$ model in Section 3.
-See the Github repo of this model from  [^1] https://github.com/pytorch/pytorch.
-As mentioned in the Experiment section 4, we adopt the pretrained PPO model from Stable baselines 3 (Raffin et al.
-2021) as the agent to be explained.
-Check the Hugging Face repositories for detailed information on the PPO agents for Enduro  [^2] https://huggingface.co/sb3/ppo-EnduroNoFrameskip-v4, Seaquest  [^3] https://huggingface.co/sb3/ppo-SeaquestNoFrameskip-v4, and MsPacman  [^4] https://huggingface.co/sb3/ppo-MsPacmanNoFrameskip-v4.
-Training details
-We conduct the training process on a server with an A40 GPU, using 50 epochs for SMB, 36 epochs for Enduro, 52 epochs for Seaquest, and 136 epochs for Ms.
-Pac-Man.
-The optimization is done using Adam.
-See Table 6 for hyperparameters.
-Hyperparameter value
-Learning rate 0.00001
-Training batch size 16
-Testing batch size 1
-Validating batch size 16
-$\lambda_{avg}$ 0.3
-$\lambda_{e}$ 1.0
-$\lambda_{\text{smooth}}$ 1.0
-$\lambda_{L2}$ 0.01
-Table 6: Hyperparameters
-Three random seeds used for training: 42, 13, 62.
-Appendix H Code
-Code can be found at: https://anonymous.4open.science/r/I6eV3cksE5AgZ7-EB88
-Appendix I Counterfactual explanations of Super Mario Bros environment
-In this section, we show more counterfactual explanations on the Super Mario Bros(Kauten 2018) environment.
-In Table 7 below, each row represents a random state from the environment.
-The original state column shows the state, the VisionMask column displays the importance map of the target action (label), and the Counterfactuals column presents the corresponding counterfactuals constructed by removing the most important regions indicated by the importance map.
-We detail the counterfactual explanations in Table 7 below:
-  • In the first row, Mario encounters the monster Goomba and two pipes, with the bush and Goomba receiving an importance score of 31.62%, while the right pipe receives 29.2%.
-Counterfactual scenarios involving the removal of the Goomba and the top section of the second green pipe both result in a change of action from accelerating while jumping right to simply jumping, which may indicate that those regions are the cause of action 4 (jump right with acceleration).
-  • In the second row, Mario is positioned between two brown bricks, with the left brick receiving an importance score of 25.23%, the right brick receiving 29.15%, and Mario and the bush each receiving 25.23%.
-It’s worth noting that jumping is necessary in the original state for Mario to pass these obstacles, but it becomes unnecessary in the counterfactual scenario.
-This observation suggests that the bricks may be the cause of Action 4, as when they are removed, jumping is no longer necessary.
-  • The third row is identical to the example presented in the counterfactual explanation section of the Super Mario Bros (SMB) environment in the paper.
-  • In the fourth column, Mario encounters a brick wall that only requires a jump (without acceleration) to pass through, with the brick wall receiving an importance score of 28.78%.
-The first counterfactual, removing the brick, indicates a change in action from Action 2 (right jump) to Action 4.
-Knowing that a right jump with acceleration (Action 4) is necessary to safely jump over the gap and land on the next platform, one can easily identify the brick wall as the cause of the action.
-In addition, the observation that the second counterfactual of removing the ground behind Mario did not change the action, which in turn supports our previous explanation.
-Note that all the object importance percentages are calculated following E.
+Appendix F Background reference values used across different environments
+In Figure 5, we list the reference values we used in each environment.
+Specifically, for SMB and Enduro, we use the background color of the content in each state as the reference value.
+For Seaquest environment, since all the states share the same background, we use a single reference value for each state, which combines the empty sea background and the oxygen bar background.
+For MsPacman, we simply used a constant value of the background color blue as the reference value.
+For VizDoom, we use black background; And we use the default grey background for Highway-env
+Appendix G Radar map for quantitative results
+Figure 6: Radar map of VisionMask and five baselines in the six environments.
+We present the radar map in Figure 6 to better visualize the quantitative results.
+Each metric was averaged across all environments.
+For deletion (Del.) and LLE, we calculated the complement by subtracting the value from 100%.
+VisionMask demonstrates better faithfulness (shown in red) compared to other baselines.
+Figure 7: Additional qualitative examples compared with five baselines in three environments.
+(a) Human explanation: ”Mario jumps to the right to get the Koopa on the ground.” VisionMask correctly identified the Koopa.
+Counterfactual analysis shows that removing the two Koopas on the ground changes the action from ’Jump right’ to ’Right.’ Note that ’Right’ cannot reach the Koopa.
+(b) Human explanation: ”Mario jumps to the right to land on the platform.” VisionMask correctly identified the platform.
+Counterfactual analysis shows that removing the platform changes the action from ’Jump right’ to ’Jump.’ (c) Human explanation: ”The agent moves to the right while maintaining its current speed because it is blocked by the car in front.” VisionMask correctly identified the agent’s current location, the front car, and the track.
+Counterfactual analysis shows that removing the front cars allows the agent to accelerate instead of choosing ’Right + accelerate.’ (d) Human explanation: ”The agent moves to the right to avoid a crash with the car on the left.” VisionMask correctly identified the agent’s current location and the two cars on the left.
+Counterfactual analysis shows that removing the cars on the left enables the agent to choose ’Left + accelerate’ instead of ’Right + accelerate.’ (e) Human explanation: ”The agent moves up to rescue the diver.” VisionMask correctly identified the relative locations of the agent and the diver, as well as the oxygen bar.
+Counterfactual analysis shows that removing the diver causes the submarine to stop moving up.
+(f) Human explanation: ”The agent stays on the surface to wait for the oxygen to fully recharge.” VisionMask correctly identified the oxygen bar.
+Counterfactual analysis shows that removing the rightmost oxygen bar causes the submarine to start operating.
+Appendix H Additional qualitative explames
+We present further qualitative examples comparing VisionMask to baselines in SMB, Enduro and Seaquest environments in Figure 7.
+Appendix I Object importance percentage calculation
+Attention percentages paid to objects are calculated by summarizing the attention map on target objects.
+Existing object detection models, trained on real-world images, perform unsatisfactorily in gaming environments.
+Furthermore, most object detection models use bounding boxes to detect objects, which are not suitable for fine-grained attention percentage calculations.
+To address these challenges, we use a segmentation model as an alternative approach.
+Specifically, we employ Segment Anything (SAM) Kirillov et al.
+[2023], which automatically generates masks indicating potential target objects, denoted as $\mathbf{m}^{\text{SAM}}\in{\{0,1\}^{N^{\text{SAM}}\times{W}\times{H}}}$, where $N^{\text{SAM}}$ is the number of potential target objects.
+We then multiply the vision mask $m_{i}\in{\{0,1\}^{W\times{H}}}$ of the ground truth action $i$ element-wise with each target object mask $m^{\text{SAM}}_{i}\in{\{0,1\}^{W\times{H}}}$ to obtain attention percentages $\mathbf{p}\in{\mathbb{R}}$ for all potential target objects.
+By re-ranking $\mathbf{p}$ and filtering out values below a predefined threshold, we derive the attention percentages on target objects.
+Finally, we manually label the target objects.
+Note that we preprocess the vision mask $m_{i}$ for attention percentage calculation by setting values below the average to zero.
+This preprocessing step filters out objects that received less attention but still show high attention percentages due to their large size, such as the background sky.
+The hyper-parameters used in SAM automatic mask generator: $\text{points\_per\_side}=64,\text{pred\_iou\_thresh}=0.7,\text{stability\_score\_thresh}=0.8$.
+The threshold we set on attention percentage calculation is $0.05$.
 Table 7: Counterfactual explanations of SMB environment.
@@
 Jump right Jump right with acceleration Jump right
-Appendix J Counterfactual explanations of Atari Enduro environment
+Appendix J Code
+We have made our trained models and testing script publicly available.
+The code can be accessed at https://anonymous.4open.science/r/I6eV3cksE5AgZ7-EB88.
+The full set of code will be published in the camera-ready version.
+Appendix K Counterfactual explanations of Super Mario Bros environment
+In this section, we show more counterfactual explanations on the Super Mario BrosKauten [2018] environment.
+In Table 7 below, each row represents a random state from the environment.
+The original state column shows the state, the VisionMask column displays the importance map of the target action (label), and the Counterfactuals column presents the corresponding counterfactuals constructed by removing the most important regions indicated by the importance map.
+We detail the counterfactual explanations in Table 7 below:
+  • In the first row, Mario encounters the monster Goomba and two pipes, with the bush and Goomba receiving an importance score of 31.62%, while the right pipe receives 29.2%.
+Counterfactual scenarios involving the removal of the Goomba and the top section of the second green pipe both result in a change of action from accelerating while jumping right to simply jumping, which may indicate that those regions are the cause of action 4 (jump right with acceleration).
+  • In the second row, Mario is positioned between two brown bricks, with the left brick receiving an importance score of 25.23%, the right brick receiving 29.15%, and Mario and the bush each receiving 25.23%.
+It’s worth noting that jumping is necessary in the original state for Mario to pass these obstacles, but it becomes unnecessary in the counterfactual scenario.
+This observation suggests that the bricks may be the cause of Action 4, as when they are removed, jumping is no longer necessary.
+  • The third row is identical to the example presented in the counterfactual explanation section of the Super Mario Bros (SMB) environment in the paper.
+  • In the fourth column, Mario encounters a brick wall that only requires a jump (without acceleration) to pass through, with the brick wall receiving an importance score of 28.78%.
+The first counterfactual, removing the brick, indicates a change in action from Action 2 (right jump) to Action 4.
+Knowing that a right jump with acceleration (Action 4) is necessary to safely jump over the gap and land on the next platform, one can easily identify the brick wall as the cause of the action.
+In addition, the observation that the second counterfactual of removing the ground behind Mario did not change the action, which in turn supports our previous explanation.
+Note that all the object importance percentages are calculated following I.
+Table 8: Counterfactual explanations of Atari Enduro environment. Note that the collision happens in the original state in the third row.
+Original state VisionMask Counterfactuals
+[Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
+Action 3: left Action 8: left + accelerate
+[Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
+Action 1: accelerate Action 7: right + accelerate
+[Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
+Action 3: left Action 8: left + accelerate
+[Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
+Action 7: right + accelerate Action 8: left + accelerate
+Appendix L Counterfactual explanations of Atari Enduro environment
 In this section, we presents counterfactual explanations on the Atari Enduro environment.
@@
 It’s obvious that the right directioned race track is the main cause of action of truing right with accelerating.
-Table 8: Counterfactual explanations of Atari Enduro environment. Note that the collision happens in the original state in the third row.
+Table 9: Counterfactual explanations of Atari Seaquest environment
 Original state VisionMask Counterfactuals
 [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
-Action 3: left Action 8: left + accelerate
+Action 13: down + fire Action 8: down + right
 [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
-Action 1: accelerate Action 7: right + accelerate
+Action 10: up + fire Action 6: up + right
 [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
-Action 3: left Action 8: left + accelerate
-[Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
-Action 7: right + accelerate Action 8: left + accelerate
-Appendix K Counterfactual explanations of Atari Seaquest environment
+Action 17: down + left + fire Action 4: left
+Appendix M Counterfactual explanations of Atari Seaquest environment
 In this section, we presents counterfactual explanations on the Atari Seaquest environment.
@@
 The oxygen level can easily be explained as the main cause in that counterfactual scenario.
-Table 9: Counterfactual explanations of Atari Seaquest environment
-Original state VisionMask Counterfactuals
-[Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
-Action 13: down + fire Action 8: down + right
-[Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
-Action 10: up + fire Action 6: up + right
-[Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
-Action 17: down + left + fire Action 4: left